### v1

# OpenECAD: An Efficient Visual Language Model for Computer-Aided Design

Abstract
Computer-aided design (CAD) tools are utilized in the manufacturing industry for modeling everything from cups to spacecraft. These programs are complex to use and typically require years of training and experience to master. Structured and well-constrained 2D sketches and 3D constructions are crucial components of CAD modeling. A well-executed CAD model can be seamlessly integrated into the manufacturing process, thereby enhancing production efficiency. Deep generative models of 3D shapes and 3D object reconstruction models has garnered significant research interest. However, most of these models are represented in discrete forms. Moreover, the few models based on CAD operations often have substantial input restrictions. In this work, we fine-tuned pre-trained models to create OpenECAD (0.55B, 0.89B, and 4.2B), leveraging the visual, logical, coding, and general capabilities of visual language models. OpenECAD can process images of 3D designs as input and generate highly structured 2D sketches and 3D construction commands. These outputs can be directly used with existing CAD tools’ APIs to generate project files. To train our network, we created a new CAD dataset. This dataset is based on existing public CAD datasets, with adjustments and augmentations to meet the requirements of  VLM training.
Index Terms:
small language model, visual language model, computer aided design, automated CAD design, 2D-3D shape modeling, geometric deep learning, machine learning
I Introduction
Using 3D shape representation aligns with human intuition. In today’s digital era, computer-aided design (CAD) tools are employed in various industrial fields for 3D shape design, including automotive, aerospace, manufacturing, and architectural design. However, due to drafting conventions and the requirements for shape constraints and edit-ability, 3D shapes are still based on 2D sketches. This approach allows for the meticulous development, relation, and annotation of all design details, mirroring the precision of traditional draftsmen.
A typical 3D part drawing process involves multiple ”Sketch-Extrusion” steps. The relationship between sketch and extrusion is not strictly one-to-one; a single sketch can correspond to multiple extrusions, and one extrusion can be composed of multiple sketches. A sketch consists of multiple closed loops formed by several lines and includes constraints within or between them to ensure the sketch is fully defined. The extrusion operation generates a 3D feature based on the 2D sketch(s). Generally, there is a sequential relationship between ”Sketch-Extrusion” steps, as a sketch’s reference plane may rely on an existing face in the model, and constraints within the Sketch may depend on points or lines already present in the model. As shown in Figure 1, this is a simple process for drawing 3D parts.
Figure 1: A schematic diagram of the process for drawing 3D parts using CAD tools.
There has been extensive research on 3D model generation. However, most of this research focuses on creating computer-discretized forms of 3D shapes, such as 3D point clouds[1, 2, 3, 4, 5], voxelized shapes[6, 7, 8, 9], polygon meshes[10, 11, 12], and levelset fields[13, 14, 15, 16, 17]. This approach neglects the essence of 3D shape design—the drawing process. In practical applications, the generated models often do not fully meet the requirements, and other steps in the production process may necessitate design modifications. Consequently, modifications to the generated models are frequently needed. However, due to the absence of the drawing process, modifying these models can be less efficient than manually modeling from scratch, making these methods challenging to apply in real-world scenarios.
To address this issue, some studies have emerged with the development of neural networks. Since 3D shapes are still based on 2D sketches, these studies can be divided into 2D sketch generation and 3D CAD model generation. SketchGraphs[18] provides a dataset of constrained 2D sketches, CurveGen and TurtleGen[19] generate closed curve sketches, and CADL[20] uses language models to generate sketches. For 3D model generation, SolidGen[21] and ComplexGen[22] generate B-rep models, while Point2Cyl[23], DeepCAD[24], and Free2CAD[25] generate CAD command sequences. It is worth noting that these models often have significant input limitations. For example, they may require existing 3D point clouds, may not accept inputs to limit how to generate, or require isometric hand-drawn sketches. These constraints still differ significantly from existing manual modeling workflows, making it challenging to apply them in production processes.
Now, with the development of language models and multimodal language models, these limitations are expected to be overcome. Multimodal language models can flexibly accept various forms of input and understand their meaning, thereby constraining the output. Currently, multimodal language models already have the capability to accept model images and program code and describe their meaning in simple language. For example, GPT-4o can possess simple 3D shape’s generation and has understanding capabilities, and many language models can interpret program code. Therefore, in this work, we leverage the visual, logical, coding, and general capabilities of visual language models to make the following contributions:
  1. 1.
    Based on existing public CAD datasets, such as SketchGraphs[18], Fusion 360 Gallery[26], and DeepCAD[24], we convert these datasets into ”image-CAD operation code” pairs, including 2D sketch generation and 3D shape generation. Using currently available multimodal large language models, such as GPT-4o, we generate high-quality annotated datasets. Additionally, we utilize rendering tools to generate various forms of images to enhance the versatility of model inputs.
  2. 2.
    Based on existing public small language models and multimodal small language models, such as TinyLLaMA[27], OpenELM[28], Phi3[29], and TinyLLaVA[30], we train them using the aforementioned training sets and test their performance in CAD design.
  3. 3.
    Design a program standard for CAD operations that can interface with the APIs of various CAD tools, enabling the generated model code to be directly used in CAD applications.
II Related Work
II-A Small Language and Visual Language Models
Small language models have recently become very popular due to the growing demand for privacy and local execution. Small language models are those that are smaller in size and require fewer computational resources, allowing them to run on local devices and ensure data privacy and security. Despite their smaller size, these models can still deliver efficient performance on specific tasks, making them suitable for resource-constrained environments. Currently available small language models include Microsoft’s Phi series[29], Apple’s OpenELM[28] and TinyLLaMA[27], based on Meta’s LLaMA[31].
Multimodal small language models further expand the applications of small language models. These models can process not only text data but also understand and generate data in various modalities, such as images and audio. By integrating information from multiple modalities, these models excel in a wider range of tasks. For example, they can convert image descriptions into text or transform hand-drawn sketches into structured textual information. Currently available multimodal small language models include the TinyLLaVA series[30], based on small language models above, and Microsoft’s Phi3 Vision[29].
II-B Generative models of 3D shapes’ CAD commands
Currently, there are several models designed for generating CAD commands for 2D Sketches or 3D shapes. SketchGraphs[18] provides a dataset of constrained 2D sketches and proposes corresponding generative models. CurveGen and TurtleGen[19] aim to generate usable sketches by producing closed curves, ignoring non-essential constraint solving. CADL[20] focuses on defining sketch operations as structures similar to programming languages, leveraging language models for sketch generation.
For 3D model generation, approaches can be divided into models generating B-rep and those generating CAD commands, depending on the final output. The Boundary representation (B-rep) format is the de-facto shape representation in CAD for modeling solid and sheet objects. B-rep has poorer editability compared to other formats but can accommodate simple model modifications. SolidGen[21] and ComplexGen[22] can both generate B-rep models. Models generating CAD commands can directly output CAD command sequences. Point2Cyl[23] can convert cylindrical 3D point clouds into CAD command sequences. DeepCAD[24] provides a dataset of CAD command sequences and proposes a model capable of generating sequences randomly. Free2CAD[25] can convert hand-drawn isometric sketches into CAD command sequences.
As these are traditional small models, they have significant input and output limitations, making integration with existing manual modeling workflows challenging. However, they offer high-quality and structurally complete datasets, which provide a foundation for generating training data for multimodal small language models. To our knowledge, there are currently no multimodal small language models specifically trained for generating CAD commands for 3D shapes.
III Method
Figure 2: Overview of the OpenECAD Dataset and Model.
We now present our OpenECAD dataset and model, which includes a new CAD command sequence format designed for language models (see Figure 2). The dataset has been improved based on existing CAD datasets and utilizes large language models to generate high-quality natural language descriptions of CAD models. We fine-tuned pre-trained vision-language models based on small language models, resulting in the OpenECAD models that are compact in size and capable of running locally, specifically designed for CAD applications.
III-A Design and Generation of Datasets
CAD models can be represented in various ways. At the user interaction level with CAD software, 3D shapes are described by a series of CAD operations to create solid forms. For instance, to model a simple table, a user begins by sketching on a 2D plane, drawing a rectangle (a closed curve) and constraining its dimensions. After completing the sketch, the user extrudes it to form the tabletop and constrains its thickness. Next, the user sketches on the existing model, drawing four squares (four closed curves) and constraining their dimensions. After completing this sketch, the user extrudes these shapes to form the table legs, merging them with the previous model and constraining their height. This completes the modeling of a simple table (see Figure 3). This process constitutes a CAD operation sequence, consisting of two Sketch-Extrusion pairs. Clearly, this CAD operation sequence can be described in natural language, as demonstrated above. Additionally, the 2D view of a CAD model is the most intuitive way to express it to humans, akin to the flowchart mentioned above.
Figure 3: A CAD Command Sequence Diagram for Drawing a Table.
In this work, we aim to generate CAD operation sequences using visual language models. CAD operation sequences not only interact with CAD software but also have meanings understandable by humans, allowing modifications and applications in other design processes. To constrain the generation of CAD operation sequences, we use human-understandable inputs such as model images or textual descriptions. Therefore, our dataset includes pairs of target model images as input with CAD operation codes and descriptions as output, pairs of target models as input with CAD operation codes as output, and pairs of target model images and descriptions as input with CAD operation codes as output. These datasets encompass both 3D CAD models and pure 2D sketches to enhance the model’s ability to generate and constrain 2D sketches.
III-A1 Definition of the Code Format for CAD Operation Sequences
Mature CAD tools support a rich set of commands, most of which are simplified calls to collections of basic commands—for example, a rectangle is a collection of four line segments and some constraints. However, in practice, only a small subset of these commands is widely used. Therefore, we currently consider only the most basic and commonly used commands in our dataset (see Table I). Due to the versatility of language models, extending the dataset later is relatively easy.
TABLE I: CAD commands and their parameters.
Command Name Parameters
Curves Series
add_line start_point, end_point
add_arc start_point, end_point, mid_point
add_circle center_point, radius
Constraints Series
make_horizontal(vertical) line0
fix_size curve0, size
make_coincident point0, point1
make_parallel line0, line1
make_perpendicular line0, line1
make_tangent curve0, curve1
make_mirror curve0, curve1 (same type)
make_angle line0, line1, clockwise
Sketch and its helper
add_sketchplane origin_point, x_axis, y_axis
add_profile loops_list
add_sketch sketchplane, profile
Extrusion
add_extrude sketch, operation, type, extent_size
Sketch. Sketch is the core of CAD operations. Extrusion is built upon Sketch. Curves and constraints form closed loops within a Sketch. Multiple closed loops make up the profile of a closed region. In our representation, the definition of Sketch begins with defining the sketch reference plane, similar to most CAD tools. We then maintain a list of loops and start drawing curves and adding constraints. Once a loop is completed, it is stored in the list. After all loops are completed, the corresponding profile is generated, and finally, the sketch is created based on the reference plane and the profile. The sketch reference plane is defined by the three-dimensional coordinates of the origin and the direction of the sketch’s XY axes.
Curves and Constraints. In practice, we consider three types of widely used curve commands: drawing a line, drawing an arc, and drawing a circle. We also consider nine types of widely used constraint commands: making a line horizontal or vertical, fixing the size of a curve, making two points of curves coincident, making two lines parallel, perpendicular, or having a fixed angle, and making two curves tangent or mirrored.
Extrusion. The extrusion command extrudes a sketch from a 2D plane into a 3D body. The extrusion type can be one-sided, symmetric, or two-sided with respect to the sketch plane. The command also specifies how to merge the newly extruded 3D body with the previously created shape using one of the operations: creating a new body, or joining, cutting, or intersecting with the existing body.
All operation names are chosen to be as meaningful as possible to aid the language model in understanding and using them. With these operation commands, we describe a CAD model as a sequence of executable commands in code.
III-A2 Translation of CAD Operation Sequences
According to the definitions provided in the previous section, we can convert any CAD model covered by the described operations (in Table I) into a CAD operation sequence as defined by us. However, in practice, different CAD tools or datasets may have slight variations in their definitions of CAD operations. We need to ensure the equivalence of CAD operations before and after the conversion. In this section, we will use the three types of curves we defined as examples to illustrate (see Figure 4, from the perspective of curve degrees of freedom (DoFs), how we ensure the equivalence of CAD models or datasets in other formats when converting them to our code.
Figure 4: Comparison Diagram of Two Definitions for Line, Arc, and Circle.
Line. According to the definition, a line is defined by two points: the start point and the end point. Since each point has 2 DoFs, a line has 4 DoFs. For a line defined by a start point, direction vector, and length, the start point has 2 DoFs, the length has 1 DoF, and although the direction vector has 2 DoFs, its length information is irrelevant for the line, contributing only 1 DoF. Therefore, lines defined by both methods have 4 DoFs and can be equivalently converted between each other using an algorithm.
Arc. According to the definition, an arc is defined by three points: the start point, end point, and midpoint. Since the midpoint must lie on the perpendicular bisector of the line segment defined by the start and end points, an arc has 5 DoFs. For an arc defined by the center, absolute angles of the start and end edges, radius, and clockwise direction, the DoFs are also evidently 5. Therefore, arcs defined by both methods have 5 DoFs and can be equivalently converted between each other using an algorithm.
Circle. According to the definition, a circle is defined by its center and radius, having 3 degrees of freedom (DoFs). For a circle defined by its center and a point on the circle, there are 4 DoFs. If we ignore the orientation of the circle (since in some CAD tools, operations like mirroring and rotating use the circle’s orientation), then the point on the circle only serves to define the radius, and its specific position is internal information of the circle. Therefore, it can be reduced to 3 DoFs. Thus, circles defined by both methods have 3 DoFs and can be equivalently converted between each other using an algorithm.
Based on the above principles, we have developed scripts to convert other CAD models or datasets into CAD operation sequence codes for training our VLM.
III-A3 Generation of Natural Language Descriptions for CAD Models
Describing CAD models in natural language is crucial for understanding 3D shapes. Adding these descriptions as annotations to the CAD operation sequence code can help people quickly grasp the meaning of each code segment and swiftly locate and modify the necessary parts. To construct natural language descriptions corresponding to CAD operation sequence codes, we employed existing large language models such as GPT-4o. By using prompts and examples, we added natural language annotations to the existing CAD operation sequence codes. The prompt used was: “Please add comments to the drawing code below to indicate what shape is drawn by each section of the code. Don’t add comment for every line or arc.” To prevent the language model from adding annotations to every CAD operation (like drawing lines or arcs), we constrained the model to add annotations by operation units, such as drawing a rectangle, etc.
III-A4 Generation of CAD Model View Images
We have introduced how to define and generate CAD model codes and textual descriptions, providing text materials for VLM training. In this section, we will briefly explain how to generate image materials for VLM training, specifically by creating 2D views of CAD models. We first use the PythonOCC[32] tool to write scripts that generate step files (B-Rep format) based on the CAD operation sequence codes. Then, we render the step files using PythonOCC. Finally, we select a direction that forms a random angle with the isometric view direction vector to generate 2D views as image materials.
III-A5 Creation of CAD Dataset
Based on the above method, we generated a 3D CAD dataset derived from the DeepCAD dataset[24] and a 2D CAD dataset derived from the SketchGraphs dataset[18]. We excluded designs that were generated incorrectly or had overly long CAD command sequences due to the number of tokens that small language models can accept is limited, resulting in a total of 130,239 3D designs and 628,499 2D sketches. The statistical distribution of the number of ”Sketch-Extrusion” steps in the 3D dataset is shown in the Figure 5. The statistical distribution of the number of curves and constraints operations in the 2D dataset is shown in the Figure 6.
Figure 5: The statistical distribution of the number of ”Sketch-Extrusion” steps in OpenECAD 3D dataset.
Figure 6: The statistical distribution of the number of curves and constraints operations in OpenECAD 2D dataset.
III-B Training Vision Language Model
III-B1 The choice of language model foundation and multimodal approach
Due to the high confidentiality requirements of manufacturing enterprises and the substantial CPU and GPU resources typically occupied by CAD design tools, we chose OpenELM[28] (450M) and Phi-3 mini[29] (3.8B) with relatively small parameter sizes as the foundation for our language model. OpenELM is an efficient language model family developed by Apple. It employs a layer-wise scaling strategy to allocate parameters efficiently within each layer of the transformer model. This approach enhances accuracy while maintaining computational efficiency. Phi-3, developed by Microsoft Research, is another series of advanced small language models (SLMs). This series includes Phi-3-mini (3.8B), Phi-3-small (7B), and Phi-3-medium (14B), each with different parameter sizes. Despite their compact parameter scales, Phi-3 models achieve language understanding and reasoning capabilities comparable to larger models. These models are trained on high-quality data and optimized algorithms. It is worth noting that the context acceptance length for OpenELM (450M) is 2048 tokens. In contrast, Phi-3 has a minimum context acceptance length of 4096 tokens, with an option to use a version supporting up to 128k tokens at the cost of some content accuracy.
To endow the language model foundation with multimodal capabilities, we selected OpenAI’s CLIP[33] and Google’s [34] as visual encoders for OpenELM (450M). These models are trained on large datasets consisting of images paired with corresponding text descriptions, combining visual and textual information to make predictions.
To train the entire model with multimodal conversational abilities, we employed the LLaVA[35] method and its dataset. LLaVA combines a visual encoder and a large language model in an innovative multimodal approach, enabling comprehensive visual and language understanding. LLaVA uses GPT-4 to generate multimodal language-image instruction data. For training the model, we utilized the training framework provided by TinyLLaVA[30].
For Phi-3 mini, since Microsoft has already released its multimodal version, Phi-3 Vision (4.2B), which was trained using methods similar to LLaVA, we will directly use the pre-trained model provided by Microsoft.
III-B2 Fine-tuning on a pre-trained visual language model
To enable the general visual-language model to generate CAD code, we fine-tuned it using the LoRA[36] method on the pre-trained OpenELM-CLIP and OpenELM-SigLIP models. LoRA stands for Low-Rank Adaptation of Large Language Models. In the field of natural language processing (NLP), there is a common paradigm that involves large-scale pre-training on general domain data followed by adaptation to specific tasks or domains. However, as we pre-train larger models, fully fine-tuning all model parameters becomes less feasible due to computational costs. LoRA proposes a novel approach. Instead of full fine-tuning, LoRA freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer of the Transformer architecture. This significantly reduces the number of trainable parameters for downstream tasks.
We performed LoRA fine-tuning on the pre-trained OpenELM-CLIP (0.55B) and OpenELM-SigLIP (0.89B) models using an Nvidia GeForce GTX 1080 Ti (11 GB VRAM). The Rank and Alpha for LoRA were set to 128 and 256, respectively. Training for one epoch with our 3D dataset took approximately 60 hours. Through fine-tuning, we obtained the OpenECAD 0.55B and 0.89B models.
Similarly, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) to perform LoRA fine-tuning on the Phi-3 Vision (4.2B) model. The LoRA rank and alpha were set to 256 and 512, respectively. Training with our dataset for one epoch took approximately 20 hours. Through fine-tuning, we obtained OpenECAD 4.2B model.
IV Experiments
IV-A Evaluation Metrics
For a given 3D shape, there are often multiple methods to create it using CAD tools. Moreover, the input views of the model generally lack dimensional information, meaning the generated model only adheres to the proportional relationships visible in the image. Therefore, directly comparing the generated CAD operation sequence codes is not reasonable. To address this, we designed a scoring algorithm and created several test CAD designs (see Figure 7). We used specific views of these designs as input images for evaluation, serving as a metric for assessing the model’s generation capability.
Figure 7: Example illustrations of CAD designs for evaluating.
TABLE II: Scoring Table.
Scoring Item Score (out of 100)
Is the code executable? 10
Accuracy of 3D shape construction 90
For each Loop-Extrusion pair (out of 100)
- 2D sketch positioning 20
- Accuracy of 2D Loop 60
- 3D operation 20
IV-A1 Scoring Algorithm for Evaluating Outputs
To assess the generative capability of our model, we designed a scoring algorithm. First, the algorithm verifies if the generated code can be executed directly by the API operator to create a model. If it cannot run, it scores 0 points. If it can run, we then use a visual language model (such as GPT-4o) to compare the generated model’s view with the target input, evaluating each basic feature. Each basic feature corresponds to a 2D loop. We separately evaluate the loops in the 2D sketches during the modeling process to ensure precise assessment, as these loops may or may not be considered a single sketch during modeling.
We score the results based on executability and the correctness of the 3D shapes (see Table II). Each Loop-Extrusion pair is considered a basic 3D shape. If the number of generated Loop-Extrusion pairs is insufficient, points will be deducted proportionally. For each Loop-Extrusion pair, we score based on the reference plane position of the 2D sketch, the extrusion operation, and the drawing of the 2D loop. The specific calculation method is shown in the following formula.
$score=10R+\frac{90}{N_{s}}\sum_{i=1}^{N_{s}}(20P_{i}+20O_{i}+60L_{j})$ (1)
where $R$ denotes executability, $N_{s}$ represents the number of Loop-Extrusion pairs, $P_{i}$ and $O_{i}$ indicate the correctness of the 2D sketch position and 3D extrusion operation for the $i$-th Sketch-Extrusion pair, and $L_{j}$ denotes the accuracy of the curves within the loop. $R,P_{i},O_{i},L_{j}\in[0,1]$.
IV-B CAD Designs for Test
To test the model, we require high-quality CAD designs. We extracted approximately 100 3D designs from the training set, maintaining a ratio of 10:5:3:2 for designs containing 1, 2, 3, and 4 or more Sketch-Extrusion pairs, respectively. After filtering and removing duplicates, we obtained 57 unique 3D designs. These designs were redrawn using CAD tools such as AutoDesk Fusion, and appropriate views were selected to generate the input images. Some examples are shown in the Figure 7. The resolution of the images is 640x400.
IV-C Evaluation Results
Figure 8: Comparison of some successful rendered test results illustrations.
IV-C1 The Inference Speed of the Models
The input was standardized to a 640x400 image of a 3D shape, with the prompt:
”This image is a view of a 3D model from a certain angle. Please try to use Python-style APIs to render this model.”
For OpenECAD 0.55B and 0.89B, we used an Nvidia GeForce GTX 1080 Ti (11 GB VRAM) for inference, with a peak memory usage of less than 5GB. Getting a 3D shape’s code took approximately 20 seconds. For OpenECAD 4.2B, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) for inference, with a peak memory usage of less than 17GB. Getting a 3D shape’s code took approximately 45 seconds.
IV-C2 Generating Code and Rendering
We tested OpenECAD versions 0.55B, 0.89B, and 4.2B, and also tested GPT-4o. Since GPT-4o was not trained for the use case described in this paper, we used the following prompt for testing:
”I am going to provide you with a 3D shape’s corresponding Python code for its modeling process. After that, I will input images of other 3D shapes, and please provide the corresponding Python code for them. ””OpenECAD style Code Example.”” This image is a view of a 3D model from a certain angle. Please write the codes of it in the format above.”
During the testing of OpenECAD, the maximum number of new tokens that 0.55B, 0.89B, and 4.2B models can generate is 1536, 1024, and 3092, respectively. Due to these context limitations, there are instances where the generated code is incomplete (see Table III). GPT-4o, due to its support for a longer context, did not encounter issues with incomplete code generation.
Similar to generating datasets, we use the PythonOCC tool to write a library that attempts to directly run the generated code to create CAD projects and output them as STEP files (B-Rep format). Then, we use PythonOCC to render the STEP files to obtain views. Since the generated programs from the model cannot ensure complete accuracy, errors may occur that the API operator cannot handle, leading to model rendering failures (see Table III).
Based on the scoring method mentioned above, we obtained the scores, the number of executable results, the number of incomplete results, the number of runtime errors, and the number of completely correct results for each model, as shown in the Table III.
TABLE III: Evaluation Results Table.
Model Overall Executable Unfinished Run with Completely
Score Errors Correct
0.55B 30.77 21 21 15 11
0.89B 29.74 17 35 5 16
4.2B 40.63 37 4 16 10
GPT-4o 11.40 38 0 19 0
IV-D Analysis of OpenECAD’s outputs
Figure 9: Typical errors in generated results.
From Table III, it is evident that OpenECAD 0.55B and 0.89B perform similarly. However, the 0.89B model uses more tokens for image input, reducing the number of new tokens it can generate. This leads to a higher number of incomplete outputs compared to the 0.55B model, resulting in a slightly lower overall score. The 4.2B model has the highest relative score but generates fewer fully correct results than the 0.55B and 0.89B models. This could be because its longer context length affects its code generation capability. All OpenECAD models significantly outperform the unfinetuned GPT-4o, whose only advantage is its extremely long context length, allowing it to complete the entire code.
In the generated results, several typical errors were observed (see Figure 9), including:
  • Case 1: For complex irregular shapes, the corresponding sketch could not be generated correctly.
  • Case 2: Errors in generating the 2D sketch profile, possibly due to intersections between loops leading to damaged extrusions.
  • Case 3: Incorrect extrusion size ratios.
  • Case 4: Incorrect sketch positioning in subsequent steps when multiple ”Sketch-Extrusion” steps are present.
  • Case 5: Difficulty in distinguishing between solid and hollow cylinders, as the relevant features are hard to observe in this view.
  • Case 6: Difficulty in distinguishing between convex and concave shapes, leading to incorrect model generation.
Based on the above results and typical errors, we can infer the current shortcomings of the OpenECAD model:
  • Weak ability to handle complex sketches, unable to reasonably determine the dimensions of various curves.
  • Poor capability in generating correctly sized dimensions.
  • Limited ability to construct complex models.
  • Insufficient detail recognition when interpreting drawings.
  • Lack of compact command sequences, leading to overly long contexts.
Based on the above, our future work will focus on expanding the types of CAD operations supported in the dataset, enhancing the model’s visual resolution and segmentation capabilities, improving the language model’s coding abilities, and increasing the model’s ability to control size and proportion.
IV-E Future Applications
Figure 10: An example of a real-world application of the model.
Currently, OpenECAD primarily serves as a simple CAD operation generation model, focusing on generating models from images. However, by improving the dataset, we can train models using similar methods and apply them to downstream applications (like Figure 10). These applications include, but are not limited to:
  • Assisting users in operating CAD design tools by guiding them on how to proceed with CAD operations to build or modify models based on their requirements and existing CAD actions.
  • Enhancing the ability to handle user CAD needs, such as design reusability and designing connectors for existing parts.
  • Introducing CAE-related knowledge to the model to address some CAE issues, such as structural reinforcement, material reduction, and kinematic simulation, during the CAD phase.
V Discussion and Conclusion
For CAD model generation, our approach has some limitations. Currently, we only consider three of the most widely used curve command types (lines, arcs, and circles), but other curve commands can be easily added. We also only consider using a single image as a reference, with the model attempting to generate the complete output result rather than step-by-step outputs. Not every CAD command sequence code currently generates a topologically valid shape. Our network cannot guarantee the correctness of its output. In practice, with sufficient context length, the number of failed executions of the generated CAD command sequence codes is relatively low.
To address these limitations, more work needs to be done. Enhancements to the dataset should include support for multi-view perspectives, complex curves, and textual requirements. The code generation process should be improved to better integrate with CAD tools, allowing the model to generate code incrementally, similar to how humans draw while referencing the design, rather than outputting the entire CAD code at once.
In summary, we have introduced OpenECAD, a visual language model and its accompanying dataset designed for generating CAD operation sequence codes. OpenECAD aims to address the challenge of CAD model generation using visual language models and has successfully generated some relatively simple 3D shapes.
Acknowledgment
We extend our gratitude to DeepCAD[24] and SketchGraphs[18] for their contributions to CAD dataset development, to TinyLLaVA[30] for providing the visual language model training framework, and to Apple[28] and Microsoft[29] for developing language models that empower developers.

### v2

@@
 A well-executed CAD model can be seamlessly integrated into the manufacturing process, thereby enhancing production efficiency.
-Deep generative models of 3D shapes and 3D object reconstruction models has garnered significant research interest.
+Deep generative models of 3D shapes and 3D object reconstruction models have garnered significant research interest.
 However, most of these models are represented in discrete forms.
 Moreover, the few models based on CAD operations often have substantial input restrictions.
-In this work, we fine-tuned pre-trained models to create OpenECAD (0.55B, 0.89B, and 4.2B), leveraging the visual, logical, coding, and general capabilities of visual language models.
+In this work, we fine-tuned pre-trained models to create OpenECAD (0.55B, 0.89B, 2.4B, 3.2B and 4.2B), leveraging the visual, logical, coding, and general capabilities of visual language models.
 OpenECAD can process images of 3D designs as input and generate highly structured 2D sketches and 3D construction commands.
@@
 This dataset is based on existing public CAD datasets, with adjustments and augmentations to meet the requirements of VLM training.
-Index Terms:
+keywords:
 small language model, visual language model, computer aided design, automated CAD design, 2D-3D shape modeling, geometric deep learning, machine learning
-I Introduction
+journal: Computer-Aided Design
+[…]
+[…]
+1 Introduction
 Using 3D shape representation aligns with human intuition.
@@
 A sketch consists of multiple closed loops formed by several lines and includes constraints within or between them to ensure the sketch is fully defined.
-The extrusion operation generates a 3D feature based on the 2D sketch(s).
+The extrusion operation generates a 3D feature based on the 2D sketch(es).
 Generally, there is a sequential relationship between ”Sketch-Extrusion” steps, as a sketch’s reference plane may rely on an existing face in the model, and constraints within the Sketch may depend on points or lines already present in the model.
@@
 For example, they may require existing 3D point clouds, may not accept inputs to limit how to generate, or require isometric hand-drawn sketches.
-These constraints still differ significantly from existing manual modeling workflows, making it challenging to apply them in production processes.
+These constraints significantly differ from existing manual modeling workflows, making it challenging to apply them in production processes.
 Now, with the development of language models and multimodal language models, these limitations are expected to be overcome.
@@
   1. 1.
+    We established OpenECAD dataset.
 Based on existing public CAD datasets, such as SketchGraphs[18], Fusion 360 Gallery[26], and DeepCAD[24], we convert these datasets into ”image-CAD operation code” pairs, including 2D sketch generation and 3D shape generation.
@@
   2. 2.
-    Based on existing public small language models and multimodal small language models, such as TinyLLaMA[27], OpenELM[28], Phi3[29], and TinyLLaVA[30], we train them using the aforementioned training sets and test their performance in CAD design.
+    We trained OpenECAD model.
+Based on existing public small language models and multimodal small language models, such as TinyLLaMA[27], OpenELM[28], Gemma[29], Phi-2, Phi-3-mini[30], and TinyLLaVA[31], we train them using the aforementioned training sets and test their performance in CAD design.
   3. 3.
-    Design a program standard for CAD operations that can interface with the APIs of various CAD tools, enabling the generated model code to be directly used in CAD applications.
-II Related Work
-II-A Small Language and Visual Language Models
+    We integrated the code generated by OpenECAD into CAD tools.
+We design a program that can interface with the APIs of various CAD tools, enabling the generated model code to be directly used in CAD applications.
+2 Related Work
+2.1 Small Language and Visual Language Models
 Small language models have recently become very popular due to the growing demand for privacy and local execution.
@@
 Despite their smaller size, these models can still deliver efficient performance on specific tasks, making them suitable for resource-constrained environments.
-Currently available small language models include Microsoft’s Phi series[29], Apple’s OpenELM[28] and TinyLLaMA[27], based on Meta’s LLaMA[31].
+Currently available small language models include Microsoft’s Phi series[30], Apple’s OpenELM[28], Google’s Gemma[29] and TinyLLaMA[27], based on Meta’s LLaMA[32].
 Multimodal small language models further expand the applications of small language models.
@@
 Currently available multimodal small language models include the TinyLLaVA series[31], based on small language models above, and Microsoft’s Phi3 Vision[30].
-II-B Generative models of 3D shapes’ CAD commands
+2.2 Generative models of 3D shapes’ CAD commands
 Currently, there are several models designed for generating CAD commands for 2D Sketches or 3D shapes.
@@
 To our knowledge, there are currently no multimodal small language models specifically trained for generating CAD commands for 3D shapes.
-III Method
+3 Method overview
 Figure 2: Overview of the OpenECAD Dataset and Model.
@@
 We fine-tuned pre-trained vision-language models based on small language models, resulting in the OpenECAD models that are compact in size and capable of running locally, specifically designed for CAD applications.
-III-A Design and Generation of Datasets
+4 Generation of Datasets
+In this chapter, we encompass several key components essential for OpenECAD datasets.
+We cover the design and generation of datasets (Section 4.1), the definition of the code format for CAD operation sequences (Section 4.2), and the translation of these sequences (Section 4.3).
+Additionally, we address the generation of natural language descriptions for CAD models (Section 4.4), the production of CAD model view images (Section 4.5), and the overall creation of the CAD dataset (Section 4.6).
+4.1 Design and Generation of Datasets
 CAD models can be represented in various ways.
@@
 These datasets encompass both 3D CAD models and pure 2D sketches to enhance the model’s ability to generate and constrain 2D sketches.
-III-A1 Definition of the Code Format for CAD Operation Sequences
+4.2 Definition of the Code Format for CAD Operation Sequences
 Mature CAD tools support a rich set of commands, most of which are simplified calls to collections of basic commands—for example, a rectangle is a collection of four line segments and some constraints.
 However, in practice, only a small subset of these commands is widely used.
-Therefore, we currently consider only the most basic and commonly used commands in our dataset (see Table I).
+Therefore, we currently consider only the most basic and commonly used commands in our dataset (see Table 1).
 Due to the versatility of language models, extending the dataset later is relatively easy.
-TABLE I: CAD commands and their parameters.
 Command Name Parameters
@@
 add_circle center_point, radius
-Constraints Series
-make_horizontal(vertical) line0
-fix_size curve0, size
-make_coincident point0, point1
-make_parallel line0, line1
-make_perpendicular line0, line1
-make_tangent curve0, curve1
-make_mirror curve0, curve1 (same type)
-make_angle line0, line1, clockwise
 Sketch and its helper
@@
 add_extrude sketch, operation, type, extent_size
+Table 1: CAD commands and their parameters.
 Sketch.
@@
 The sketch reference plane is defined by the three-dimensional coordinates of the origin and the direction of the sketch’s XY axes.
-Curves and Constraints.
+Curves.
 In practice, we consider three types of widely used curve commands: drawing a line, drawing an arc, and drawing a circle.
-We also consider nine types of widely used constraint commands: making a line horizontal or vertical, fixing the size of a curve, making two points of curves coincident, making two lines parallel, perpendicular, or having a fixed angle, and making two curves tangent or mirrored.
+The line is defined by its starting and ending points, the arc is determined by its starting point, ending point, and midpoint, and the circle is specified by its center and radius.
 Extrusion.
@@
 With these operation commands, we describe a CAD model as a sequence of executable commands in code.
-III-A2 Translation of CAD Operation Sequences
-According to the definitions provided in the previous section, we can convert any CAD model covered by the described operations (in Table I) into a CAD operation sequence as defined by us.
-However, in practice, different CAD tools or datasets may have slight variations in their definitions of CAD operations.
-We need to ensure the equivalence of CAD operations before and after the conversion.
-In this section, we will use the three types of curves we defined as examples to illustrate (see Figure 4, from the perspective of curve degrees of freedom (DoFs), how we ensure the equivalence of CAD models or datasets in other formats when converting them to our code.
+The pseudo code corresponding to a pair of Sketch-Extrusion operations is shown in the Algorithm 1.
+Algorithm 1 Pseudo Code Example
+$SketchPlane$ = add_sketchplane($origin$, $normal$, $x_{axis}$, $y_{axis}$)
+$Loops$ and $Curves$ are lists.
+$Line1$ = add_line($start$, $end$)
+$Line1$ is appended to $Curves$.
+$Loop$ = add_loop($Curves$)
+$Loop$ is appended to $Loops$.
+$Profile$ = add_profile($Loops$)
+$Sketch$ = add_sketch($SketchPlane$, $profile$, $position$, $size$)
+$Extrude$ = add_extrude($Sketch$, $operation$, $type$, $extent$)
+4.3 Translation of CAD Operation Sequences
+According to the definitions provided in the previous section, we can convert any CAD model covered by the described operations (in Table 1) into a CAD operation sequence as defined by us.
+However, in practice, different CAD tools or datasets may have slight variations in their definitions of CAD operations (see Figure 4).
+Therefore, we determine and transform each curve-drawing operation according to its degrees of freedom to match the syntax of the OpenECAD dataset.
+We have developed scripts to convert other CAD models or datasets into CAD operation sequence codes for training our VLM.
 Figure 4: Comparison Diagram of Two Definitions for Line, Arc, and Circle.
-Line.
-According to the definition, a line is defined by two points: the start point and the end point.
-Since each point has 2 DoFs, a line has 4 DoFs.
-For a line defined by a start point, direction vector, and length, the start point has 2 DoFs, the length has 1 DoF, and although the direction vector has 2 DoFs, its length information is irrelevant for the line, contributing only 1 DoF.
-Therefore, lines defined by both methods have 4 DoFs and can be equivalently converted between each other using an algorithm.
-Arc.
-According to the definition, an arc is defined by three points: the start point, end point, and midpoint.
-Since the midpoint must lie on the perpendicular bisector of the line segment defined by the start and end points, an arc has 5 DoFs.
-For an arc defined by the center, absolute angles of the start and end edges, radius, and clockwise direction, the DoFs are also evidently 5.
-Therefore, arcs defined by both methods have 5 DoFs and can be equivalently converted between each other using an algorithm.
-Circle.
-According to the definition, a circle is defined by its center and radius, having 3 degrees of freedom (DoFs).
-For a circle defined by its center and a point on the circle, there are 4 DoFs.
-If we ignore the orientation of the circle (since in some CAD tools, operations like mirroring and rotating use the circle’s orientation), then the point on the circle only serves to define the radius, and its specific position is internal information of the circle.
-Therefore, it can be reduced to 3 DoFs.
-Thus, circles defined by both methods have 3 DoFs and can be equivalently converted between each other using an algorithm.
-Based on the above principles, we have developed scripts to convert other CAD models or datasets into CAD operation sequence codes for training our VLM.
-III-A3 Generation of Natural Language Descriptions for CAD Models
+4.4 Generation of Natural Language Descriptions for CAD Models
 Describing CAD models in natural language is crucial for understanding 3D shapes.
@@
 By using prompts and examples, we added natural language annotations to the existing CAD operation sequence codes.
-The prompt used was: “Please add comments to the drawing code below to indicate what shape is drawn by each section of the code.
-Don’t add comment for every line or arc.” To prevent the language model from adding annotations to every CAD operation (like drawing lines or arcs), we constrained the model to add annotations by operation units, such as drawing a rectangle, etc.
-III-A4 Generation of CAD Model View Images
+The prompt used was:
+“Please add comments to the drawing code below to indicate what shape is drawn by each section of the code. Don’t add comment for every line or arc.”
+To prevent the language model from adding annotations to every CAD operation (like drawing lines or arcs), we constrained the model to add annotations by operation units, such as drawing a rectangle, etc.
+4.5 Generation of CAD Model View Images
 We have introduced how to define and generate CAD model codes and textual descriptions, providing text materials for VLM training.
@@
 Finally, we select a direction that forms a random angle with the isometric view direction vector to generate 2D views as image materials.
-III-A5 Creation of CAD Dataset
-Based on the above method, we generated a 3D CAD dataset derived from the DeepCAD dataset[24] and a 2D CAD dataset derived from the SketchGraphs dataset[18].
-We excluded designs that were generated incorrectly or had overly long CAD command sequences due to the number of tokens that small language models can accept is limited, resulting in a total of 130,239 3D designs and 628,499 2D sketches.
+4.6 Creation of CAD Dataset
+Based on the above method, we generated a 3D CAD dataset derived from DeepCAD dataset[24] and Fusion 360 Gallery Reconstruction Dataset[26].
+We excluded designs that were generated incorrectly or had overly long CAD command sequences due to the number of tokens that small language models can accept is limited, resulting in a total of 130,239 3D designs.
 The statistical distribution of the number of ”Sketch-Extrusion” steps in the 3D dataset is shown in the Figure 5.
-The statistical distribution of the number of curves and constraints operations in the 2D dataset is shown in the Figure 6.
 Figure 5: The statistical distribution of the number of ”Sketch-Extrusion” steps in OpenECAD 3D dataset.
-Figure 6: The statistical distribution of the number of curves and constraints operations in OpenECAD 2D dataset.
-III-B Training Vision Language Model
-III-B1 The choice of language model foundation and multimodal approach
-Due to the high confidentiality requirements of manufacturing enterprises and the substantial CPU and GPU resources typically occupied by CAD design tools, we chose OpenELM[28] (450M) and Phi-3 mini[29] (3.8B) with relatively small parameter sizes as the foundation for our language model.
+5 Training and Fine-tuning Vision Language Models
+In this chapter, we introduce the process of obtaining OpenECAD models.
+We discuss the choice of language models and multimodal approach (Section 5.1), as well as the fine-tuning of a pre-trained visual language model (Section 5.2).
+5.1 The choice of language models and multimodal approach
+Due to the high confidentiality requirements of manufacturing enterprises and the substantial CPU and GPU resources typically occupied by CAD design tools, we chose OpenELM[28] (450M), Gemma[29] (2B), Phi-2 (2.7B) and Phi-3 mini[30] (3.8B) with relatively small parameter sizes as the foundation for our language model.
 OpenELM is an efficient language model family developed by Apple.
@@
 This approach enhances accuracy while maintaining computational efficiency.
-Phi-3, developed by Microsoft Research, is another series of advanced small language models (SLMs).
-This series includes Phi-3-mini (3.8B), Phi-3-small (7B), and Phi-3-medium (14B), each with different parameter sizes.
+Gemma is a family of lightweight, text-to-text, decoder-only large language models developed by Google.
+Phi, developed by Microsoft Research, is another series of advanced small language models (SLMs).
+This series includes Phi-2 (2.7B), Phi-3-mini (3.8B), Phi-3-small (7B), and Phi-3-medium (14B), each with different parameter sizes.
 Despite their compact parameter scales, Phi-3 models achieve language understanding and reasoning capabilities comparable to larger models.
@@
 It is worth noting that the context acceptance length for OpenELM (450M) is 2048 tokens.
-In contrast, Phi-3 has a minimum context acceptance length of 4096 tokens, with an option to use a version supporting up to 128k tokens at the cost of some content accuracy.
-To endow the language model foundation with multimodal capabilities, we selected OpenAI’s CLIP[33] and Google’s [34] as visual encoders for OpenELM (450M).
+In contrast, Gemmma and Phi has a minimum context acceptance length of 3072 tokens, with an option to use a version supporting up to 128k tokens at the cost of some content accuracy.
+To endow the language model foundation with multimodal capabilities, we selected OpenAI’s CLIP[34] and Google’s SigLIP [35] as visual encoders for OpenELM (450M), Gemma (2B) and Phi-2 (2.7B).
 These models are trained on large datasets consisting of images paired with corresponding text descriptions, combining visual and textual information to make predictions.
@@
 For Phi-3 mini, since Microsoft has already released its multimodal version, Phi-3 Vision (4.2B), which was trained using methods similar to LLaVA, we will directly use the pre-trained model provided by Microsoft.
-III-B2 Fine-tuning on a pre-trained visual language model
-To enable the general visual-language model to generate CAD code, we fine-tuned it using the LoRA[36] method on the pre-trained OpenELM-CLIP and OpenELM-SigLIP models.
+5.2 Fine-tuning on a pre-trained visual language model
+OpenECAD 0.55B & 0.89B
+CPU Intel Core i9-7960X
+GPU Nvidia GeForce GTX 1080 Ti
+RAM DDR4-2666 32 GB
+OS Ubuntu Server 22.04 LTS
+OpenECAD 2.4B & 3.1B
+CPU AMD EPYC 9754 Processor (18 vCPU)
+GPU 2 x Nvidia GeForce RTX 4090D
+RAM 60 GB
+OS Ubuntu Server 22.04 LTS
+OpenECAD 4.2B
+CPU AMD EPYC 9754 Processor (18 vCPU)
+GPU Nvidia GeForce RTX 4090D
+RAM 60 GB
+OS Ubuntu Server 22.04 LTS
+Table 2: Development and running environment.
+To enable the general visual-language model to generate CAD code, we fine-tuned it using the LoRA[37] method on the pre-trained OpenELM-CLIP, OpenELM-SigLIP, Gemma-SigLIP and Phi-2-SigLIP models.
 LoRA stands for Low-Rank Adaptation of Large Language Models.
@@
 The Rank and Alpha for LoRA were set to 128 and 256, respectively.
+The learning rate is $10^{-4}$.
 Training for one epoch with our 3D dataset took approximately 60 hours.
 Through fine-tuning, we obtained the OpenECAD 0.55B and 0.89B models.
-Similarly, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) to perform LoRA fine-tuning on the Phi-3 Vision (4.2B) model.
+Similarly, we used 2 Nvidia GeForce RTX 4090 D (24 GB VRAM) to perform LoRA fine-tuning on the Gemma-SigLIP (2.4B) and Phi-2-SigLIP (3.1B) model.
+The Rank and Alpha for LoRA were set to 128 and 256, respectively.
+The learning rate is $10^{-4}$.
+Training for one epoch with our 3D dataset took approximately 14 hours.
+Through fine-tuning, we obtained the OpenECAD 2.4B and 3.1B models.
+We used an Nvidia GeForce RTX 4090 D (24 GB VRAM) to perform LoRA fine-tuning on the Phi-3 Vision (4.2B) model.
 The LoRA rank and alpha were set to 256 and 512, respectively.
+The learning rate is $10^{-4}$.
 Training with our dataset for one epoch took approximately 20 hours.
 Through fine-tuning, we obtained OpenECAD 4.2B model.
-IV Experiments
-IV-A Evaluation Metrics
+The specific development, training, and runtime environments are shown in the Table 2.
+The loss curves for the OpenECAD 0.55B, 0.89B, 2.4B, 3.2B and 4.2B models during training are illustrated in the Figure 6.
+Due to the different batch sizes used during training, the loss curves are proportionally scaled.
+As shown, the models converge after a single round of fine-tuning.
+Figure 6: The loss curves for the OpenECAD 0.55B, 0.89B, 2.4B, 3.2B and 4.2B models during training.
+6 Experiments
+In this chapter, we evaluate the ability of OpenECAD models to generate CAD designs based on input images and analyze the results.
+As there are no existing CAD models that use images for CAD generation, we first define evaluation metrics (Section 6.1) and create CAD designs for testing (Section 6.2).
+Following this, we test the models, obtain the results (Section 6.3), and analyze them (Section 6.4).
+6.1 Evaluation Metrics
 For a given 3D shape, there are often multiple methods to create it using CAD tools.
@@
 Figure 7: Example illustrations of CAD designs for evaluating.
-TABLE II: Scoring Table.
 Scoring Item Score (out of 100)
@@
 - 3D operation 20
-IV-A1 Scoring Algorithm for Evaluating Outputs
+Table 3: Scoring Table.
+6.1.1 Scoring Algorithm for Evaluating Outputs
 To assess the generative capability of our model, we designed a scoring algorithm.
@@
 We separately evaluate the loops in the 2D sketches during the modeling process to ensure precise assessment, as these loops may or may not be considered a single sketch during modeling.
-We score the results based on executability and the correctness of the 3D shapes (see Table II).
+We score the results based on executability and the correctness of the 3D shapes (see Table 3).
 Each Loop-Extrusion pair is considered a basic 3D shape.
@@
 $R,P_{i},O_{i},L_{j}\in[0,1]$.
-IV-B CAD Designs for Test
+6.2 CAD Designs for Test
 To test the model, we require high-quality CAD designs.
@@
 The resolution of the images is 640x400.
-IV-C Evaluation Results
+6.3 Evaluation Results
 Figure 8: Comparison of some successful rendered test results illustrations.
-IV-C1 The Inference Speed of the Models
+6.3.1 The Inference Speed of the Models
 The input was standardized to a 640x400 image of a 3D shape, with the prompt:
@@
 Getting a 3D shape’s code took approximately 20 seconds.
-For OpenECAD 4.2B, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) for inference, with a peak memory usage of less than 17GB.
+For OpenECAD 2.4B, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) for inference, with a peak memory usage of less than 9GB.
+For OpenECAD 3.2B and 4.2B, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) for inference, with a peak memory usage of less than 18GB.
 Getting a 3D shape’s code took approximately 45 seconds.
-IV-C2 Generating Code and Rendering
-We tested OpenECAD versions 0.55B, 0.89B, and 4.2B, and also tested GPT-4o. Since GPT-4o was not trained for the use case described in this paper, we used the following prompt for testing:
+6.3.2 Generating Code and Rendering
+We tested OpenECAD versions 0.55B, 0.89B, 2.4B, 3.2B and 4.2B, and also tested GPT-4o.
+Some successfully rendered test results are illustrated in Figure 8.
+Since GPT-4o was not trained specifically for the use case described in this paper, we used the following prompt for testing:
 ”I am going to provide you with a 3D shape’s corresponding Python code for its modeling process.
 After that, I will input images of other 3D shapes, and please provide the corresponding Python code for them.
-””OpenECAD style Code Example.”” This image is a view of a 3D model from a certain angle.
+$<$OpenECAD style Code Example$>$ This image is a view of a 3D model from a certain angle.
 Please write the codes of it in the format above.”
-During the testing of OpenECAD, the maximum number of new tokens that 0.55B, 0.89B, and 4.2B models can generate is 1536, 1024, and 3092, respectively.
-Due to these context limitations, there are instances where the generated code is incomplete (see Table III).
+During the testing of OpenECAD, the maximum number of new tokens that 0.55B, 0.89B, 3.2B and 4.2B models can generate is 1536, 1024, 2048, 3072 and 3072, respectively.
+Due to these context limitations, there are instances where the generated code is incomplete.
 GPT-4o, due to its support for a longer context, did not encounter issues with incomplete code generation.
@@
 Then, we use PythonOCC to render the STEP files to obtain views.
-Since the generated programs from the model cannot ensure complete accuracy, errors may occur that the API operator cannot handle, leading to model rendering failures (see Table III).
-Based on the scoring method mentioned above, we obtained the scores, the number of executable results, the number of incomplete results, the number of runtime errors, and the number of completely correct results for each model, as shown in the Table III.
-TABLE III: Evaluation Results Table.
+Since the generated programs from the model cannot ensure complete accuracy, errors may occur that the API operator cannot handle, leading to model rendering failures.
+Based on the scoring method mentioned above, we obtained the scores, the number of executable results, the number of incomplete results, the number of runtime errors, and the number of completely correct results for each model, as shown in the Table 4.
 Model Overall Executable Unfinished Run with Completely
 Score Errors Correct
-0.55B 30.77 21 21 15 11
-0.89B 29.74 17 35 5 16
-4.2B 40.63 37 4 16 10
-GPT-4o 11.40 38 0 19 0
-IV-D Analysis of OpenECAD’s outputs
+0.55B 46.52 36 21 0 15
+0.89B 37.19 22 35 0 19
+2.4B 59.86 50 7 0 20
+3.1B 55.39 53 0 4 16
+4.2B 51.91 51 4 2 14
+GPT-4o 15.44 52 0 5 4
+Table 4: Evaluation Results Table.
+As shown in Table 4, OpenECAD 2.4B achieved the highest overall score and the most fully correct outputs.
+Both 0.55B and 0.89B models produced a significant amount of incomplete code.
+The 0.89B model’s higher token consumption for image inputs reduced the number of new tokens it could generate, leading to more incomplete outputs compared to the 0.55B model, and thus a lower overall score.
+The 3.1B and 4.2B models achieved relatively high scores, but the number of fully correct generations was lower than that of the 0.55B and 0.89B models.
+All OpenECAD models demonstrated significant advantages over the unrefined GPT-4o, whose sole advantage was its ability to handle the entire code due to its extended context length.
+6.4 Analysis of OpenECAD’s outputs
+6.4.1 Code Generation Analysis
+For the generated code, the language model effectively adheres to the OpenECAD dataset’s syntax, even without being explicitly instructed on the specific details of the syntax.
+Regarding the curves in the sketch, the language model successfully ensures that consecutive curves are connected end-to-end, with the first and last curves also joined, thus maintaining closure.
+Additionally, the language model adeptly manages the positions of points to ensure the profile’s validity.
+For the fully generated code, the vast majority can be executed and rendered directly.
+For generated code with runtime errors, this includes both execution crashes and issues where the B-Rep file fails to render correctly.
+6.4.2 Generation Error Analysis
 Figure 9: Typical errors in generated results.
-From Table III, it is evident that OpenECAD 0.55B and 0.89B perform similarly.
-However, the 0.89B model uses more tokens for image input, reducing the number of new tokens it can generate.
-This leads to a higher number of incomplete outputs compared to the 0.55B model, resulting in a slightly lower overall score.
-The 4.2B model has the highest relative score but generates fewer fully correct results than the 0.55B and 0.89B models.
-This could be because its longer context length affects its code generation capability.
-All OpenECAD models significantly outperform the unfinetuned GPT-4o, whose only advantage is its extremely long context length, allowing it to complete the entire code.
 In the generated results, several typical errors were observed (see Figure 9), including:
-  • Case 1: For complex irregular shapes, the corresponding sketch could not be generated correctly.
-  • Case 2: Errors in generating the 2D sketch profile, possibly due to intersections between loops leading to damaged extrusions.
-  • Case 3: Incorrect extrusion size ratios.
-  • Case 4: Incorrect sketch positioning in subsequent steps when multiple ”Sketch-Extrusion” steps are present.
-  • Case 5: Difficulty in distinguishing between solid and hollow cylinders, as the relevant features are hard to observe in this view.
-  • Case 6: Difficulty in distinguishing between convex and concave shapes, leading to incorrect model generation.
+  • 1.
+    Case 1 and Case 2: For complex and irregular shapes, the corresponding sketches often fail to generate correctly, or the profiles of the 2D sketches are inaccurately created.
+Possible reasons include the vision tower’s low resolution, which hinders accurate recognition of complex shapes, and the language model’s inability to comprehend complex shapes from the CAD operation code.
+These issues ultimately result in incorrect loops or intersecting loops, leading to erroneous sketch profiles and broken extrusions.
+  • 2.
+    Case 3 and Case 7: The model lacks spatial reasoning capabilities, resulting in incorrect handling of the size proportions between sketches and extrusions.
+This leads to extrusions that are either too long or too small, causing discrepancies in size relative to other parts.
+  • 3.
+    Case 4: When multiple ”Sketch-Extrusion” steps are involved, the positions of the sketches in subsequent steps are often incorrect, leading to errors in the final overall model.
+  • 4.
+    Case 5: Due to the minimal color difference between the outer surface of a solid cylinder and the inner wall of a hollow cylinder, the primary distinguishing feature is the inner wall circle at the top.
+This makes it difficult for the model to differentiate between solid and hollow cylinders, leading to confusion.
+This issue also extends to other similar shapes with holes.
+  • 5.
+    Case 6: The difficulty in distinguishing concave and convex shapes results in erroneous models.
+This issue may arise because the vision tower can detect these shapes but cannot differentiate between them, or because the language model fails to generate the correct code, leading to sketch errors.
+  • 6.
+    Case 8: Missing part(s). For some relatively subtle features, the model struggles to accurately identify them, resulting in missing features.
 Based on the above results and typical errors, we can infer the current shortcomings of the OpenECAD model:
-  • Weak ability to handle complex sketches, unable to reasonably determine the dimensions of various curves.
-  • Poor capability in generating correctly sized dimensions.
-  • Limited ability to construct complex models.
-  • Insufficient detail recognition when interpreting drawings.
-  • Lack of compact command sequences, leading to overly long contexts.
-Based on the above, our future work will focus on expanding the types of CAD operations supported in the dataset, enhancing the model’s visual resolution and segmentation capabilities, improving the language model’s coding abilities, and increasing the model’s ability to control size and proportion.
-IV-E Future Applications
+  • 1.
+    The model has weak handling capabilities for complex sketches.
+This includes issues such as the vision tower’s low resolution, which leads to an inability to recognize shapes, and the language model’s failure to correctly process complex shapes, resulting in erroneous sketch generation.
+  • 2.
+    The generated models have poor size accuracy, particularly in handling the proportions between sketch dimensions and extrusion sizes.
+This results in mismatched parts, as the model fails to ensure consistent sizes between the sketches and extrusions of different components.
+  • 3.
+    For complex shapes involving multiple ”Sketch-Extrusion” steps, the model struggles to generate accurate results.
+It has difficulty remembering the previously drawn parts through the context, leading to errors.
+  • 4.
+    Due to the current work focusing only on basic instructions without a combined instruction set like drawing rectangles in one command, the CAD operation codes for 3D shapes tend to be lengthy.
+Moreover, due to contextual limitations of the model, especially with OpenECAD 0.55B and 0.89B using OpenELM which supports only 2048 tokens, and considering that image inputs also consume tokens, there are situations where the code cannot be completed.
+6.5 Future Work
+6.5.1 Refinement
+Based on the above analysis, our future work will focus on the following areas:
+  • 1.
+    Enhancing the dataset involves several key strategies: expanding support for additional CAD instructions, annotating critical shape dimensions in input images to aid accurate size-based shape drawing by the model, selecting optimal angles for input images to showcase all features comprehensively, and employing diverse rendering tools to enhance the model’s ability to recognize shapes under varied lighting conditions and artistic styles.
+  • 2.
+    Improving the visual model is crucial as the current resolution limits its ability to distinguish between certain shapes.
+Therefore, training a visual model capable of handling higher resolutions is necessary.
+Additionally, integrating an image segmentation module like FastSAM[38] before the visual model can aid in identifying which parts compose the 3D shapes.
+  • 3.
+    Improving the language model’s capabilities is essential, as the current model supports a limited context length and has weak code generation abilities, resulting in incomplete or erroneous code.
+We can address this by using larger language models that support longer contexts, such as Phi-3-medium or LLaMA-3.
+Additionally, implementing Retrieval-Augmented Generation (RAG) [39] and prompting techniques can further enhance the model’s contextual understanding and code generation capabilities.
+  • 4.
+    Enhancing feedback integration with CAD tools.
+By adopting a conversational approach, partly generated CAD instructions can be promptly sent to the CAD tool for rendering.
+The rendering results can then be immediately returned to the model, helping it understand the current drawing status, identify potential errors, correct them, and continue generating code based on the target shape and the already rendered results.
+At the same time, we have added a Constraints section to the dataset.
+We consider nine types of widely used constraint commands (see Table 5): making a line horizontal or vertical, fixing the size of a curve, making two points of curves coincident, making two lines parallel, perpendicular, or at a fixed angle, and making two curves tangent or mirrored.
+These constraints reduce the degrees of freedom in the sketch.
+A sketch with exactly zero degrees of freedom is fully defined, while one with fewer than zero degrees of freedom is over-constrained.
+We generated a 2D CAD dataset derived from the SketchGraphs dataset[18], resulting in a total of 628,499 2D sketches.
+The statistical distribution of the number of curves and constraints operations in the 2D dataset is shown in the Figure 10.
+Command Name Parameters
+Constraints Series
+make_horizontal(vertical) line0
+fix_size curve0, size
+make_coincident point0, point1
+make_parallel line0, line1
+make_perpendicular line0, line1
+make_tangent curve0, curve1
+make_mirror curve0, curve1 (same type)
+make_angle line0, line1, clockwise
+Table 5: CAD commands and their parameters.
+Figure 10: The statistical distribution of the number of curves and constraints operations in OpenECAD 2D dataset.
+6.5.2 Applications
 Figure 11: An example of a real-world application of the model.
@@
 These applications include, but are not limited to:
+  • 1.
     Assisting users in operating CAD design tools by guiding them on how to proceed with CAD operations to build or modify models based on their requirements and existing CAD actions.
+  • 2.
     Enhancing the ability to handle user CAD needs, such as design reusability and designing connectors for existing parts.
+  • 3.
     Introducing CAE-related knowledge to the model to address some CAE issues, such as structural reinforcement, material reduction, and kinematic simulation, during the CAD phase.
-V Discussion and Conclusion
+7 Discussion and Conclusion
 For CAD model generation, our approach has some limitations.
@@
 Our network cannot guarantee the correctness of its output.
-In practice, with sufficient context length, the number of failed executions of the generated CAD command sequence codes is relatively low.
+In practice, when the context length is sufficient, the number of execution failures in the generated CAD command sequence code is relatively low.
+However, when the context length is insufficient, it becomes challenging to generate the code completely.
 To address these limitations, more work needs to be done.
@@
 OpenECAD aims to address the challenge of CAD model generation using visual language models and has successfully generated some relatively simple 3D shapes.
+We offer 5 sizes of OpenECAD models, ranging from 0.55B to 4.2B parameters.
+CRediT authorship contribution statement
+Zhe Yuan: Software, Writing – original draft, Writing – review & editing, Project administation. Jianqi Shi: Writing – review & editing, Supervision. Yanhong Huang: Supervision.
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Data availability
+Data will be made available on request.
+Declaration of generative AI and AI-assisted technologies in the writing process
+During the preparation of this work the author(s) used ChatGPT in order to correct the possible wrong expression in English.
+After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
 Acknowledgment
-We extend our gratitude to DeepCAD[24] and SketchGraphs[18] for their contributions to CAD dataset development, to TinyLLaVA[30] for providing the visual language model training framework, and to Apple[28] and Microsoft[29] for developing language models that empower developers.
+This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.

### v3

@@
-OpenECAD: An Efficient Visual Language Model for Computer-Aided Design
+OpenECAD: An Efficient Visual Language Model for Editable 3D-CAD Design
 Abstract
@@
 Deep generative models of 3D shapes and 3D object reconstruction models have garnered significant research interest.
-However, most of these models are represented in discrete forms.
+However, most of these models produce discrete forms of 3D objects that are not editable.
 Moreover, the few models based on CAD operations often have substantial input restrictions.
-In this work, we fine-tuned pre-trained models to create OpenECAD (0.55B, 0.89B, 2.4B, 3.2B and 4.2B), leveraging the visual, logical, coding, and general capabilities of visual language models.
-OpenECAD can process images of 3D designs as input and generate highly structured 2D sketches and 3D construction commands.
+In this work, we fine-tuned pre-trained models to create OpenECAD models (0.55B, 0.89B, 2.4B and 3.1B), leveraging the visual, logical, coding, and general capabilities of visual language models.
+OpenECAD models can process images of 3D designs as input and generate highly structured 2D sketches and 3D construction commands, ensuring that the designs are editable.
 These outputs can be directly used with existing CAD tools’ APIs to generate project files.
-To train our network, we created a new CAD dataset.
-This dataset is based on existing public CAD datasets, with adjustments and augmentations to meet the requirements of VLM training.
+To train our network, we created a series of OpenECAD datasets.
+These datasets are derived from existing public CAD datasets, adjusted and augmented to meet the specific requirements of vision language model (VLM) training.
+Additionally, we have introduced an approach that utilizes dependency relationships to define and generate sketches, further enriching the content and functionality of the datasets.
 keywords:
-small language model, visual language model, computer aided design, automated CAD design, 2D-3D shape modeling, geometric deep learning, machine learning
-journal: Computer-Aided Design
+small language model; visual language model; computer aided design; geometric deep learning
 […]
@@
 1 Introduction
-Using 3D shape representation aligns with human intuition.
 In today’s digital era, computer-aided design (CAD) tools are employed in various industrial fields for 3D shape design, including automotive, aerospace, manufacturing, and architectural design.
@@
 This approach allows for the meticulous development, relation, and annotation of all design details, mirroring the precision of traditional draftsmen.
-A typical 3D part drawing process involves multiple ”Sketch-Extrusion” steps.
-The relationship between sketch and extrusion is not strictly one-to-one; a single sketch can correspond to multiple extrusions, and one extrusion can be composed of multiple sketches.
+A typical 3D part drawing process involves multiple “Sketch-Extrusion” steps, where one sketch can correspond to multiple extrusions.
 A sketch consists of multiple closed loops formed by several lines and includes constraints within or between them to ensure the sketch is fully defined.
 The extrusion operation generates a 3D feature based on the 2D sketch(es).
-Generally, there is a sequential relationship between ”Sketch-Extrusion” steps, as a sketch’s reference plane may rely on an existing face in the model, and constraints within the Sketch may depend on points or lines already present in the model.
-As shown in Figure 1, this is a simple process for drawing 3D parts.
-Figure 1: A schematic diagram of the process for drawing 3D parts using CAD tools.
+Generally, there is a sequential relationship between “Sketch-Extrusion” steps, as a sketch’s reference plane may rely on an existing face in the model, and constraints within the sketch may depend on points or lines already present in the model.
 There has been extensive research on 3D model generation.
-However, most of this research focuses on creating computer-discretized forms of 3D shapes, such as 3D point clouds[1, 2, 3, 4, 5], voxelized shapes[6, 7, 8, 9], polygon meshes[10, 11, 12], and levelset fields[13, 14, 15, 16, 17].
-This approach neglects the essence of 3D shape design—the drawing process.
-In practical applications, the generated models often do not fully meet the requirements, and other steps in the production process may necessitate design modifications.
-Consequently, modifications to the generated models are frequently needed.
-However, due to the absence of the drawing process, modifying these models can be less efficient than manually modeling from scratch, making these methods challenging to apply in real-world scenarios.
-To address this issue, some studies have emerged with the development of neural networks.
-Since 3D shapes are still based on 2D sketches, these studies can be divided into 2D sketch generation and 3D CAD model generation.
-SketchGraphs[18] provides a dataset of constrained 2D sketches, CurveGen and TurtleGen[19] generate closed curve sketches, and CADL[20] uses language models to generate sketches.
-For 3D model generation, SolidGen[21] and ComplexGen[22] generate B-rep models, while Point2Cyl[23], DeepCAD[24], and Free2CAD[25] generate CAD command sequences.
-It is worth noting that these models often have significant input limitations.
-For example, they may require existing 3D point clouds, may not accept inputs to limit how to generate, or require isometric hand-drawn sketches.
-These constraints significantly differ from existing manual modeling workflows, making it challenging to apply them in production processes.
-Now, with the development of language models and multimodal language models, these limitations are expected to be overcome.
+Most of this research focuses on creating computer-discretized forms of 3D shapes, such as 3D point clouds[1, 2, 3, 4, 5], voxelized shapes[6, 7, 8, 9], polygon meshes[10, 11, 12], and levelset fields[13, 14, 15, 16, 17].
+This approach focuses on the application of precise 3D point cloud models in cyberspace, neglecting the essence of 3D shape design—the drawing process, and therefore is not suitable for the design stage of 3D models.
+With the development of neural networks, some studies have emerged focusing on generating 3D shapes based on points and surfaces.
+SolidGen[18] and ComplexGen[19] generate B-rep models, while Point2Cyl[20], DeepCAD[21], and Free2CAD[22] generate CAD command sequences.
+In practical design workflows, generated 3D models often do not fully meet design requirements and need modifications due to production processes like draft angles and forging.
+Without the drawing process, these modifications can be less efficient than manually modeling from scratch, making these methods challenging to apply.
+Additionally, these models often have input limitations, such as requiring existing 3D point clouds or detailed hand-drawn information in isometric views.
+Now, with the development of language models and multimodal language models, these limitations are being overcome.
 Multimodal language models can flexibly accept various forms of input and understand their meaning, thereby constraining the output.
-Currently, multimodal language models already have the capability to accept model images and program code and describe their meaning in simple language.
-For example, GPT-4o can possess simple 3D shape’s generation and has understanding capabilities, and many language models can interpret program code.
-Therefore, in this work, we leverage the visual, logical, coding, and general capabilities of visual language models to make the following contributions:
+Currently, multimodal language models already have the basic ability to accept 2D views or language descriptions of 3D shapes and use CAD operations to draw them.
+For example, GPT-4o can possess simple 3D shape generation and understanding capabilities, CAD-LLM[23] can manipulate engineering sketches by fine-tuning pre-trained language models to generate parametric CAD.
+Figure 1: Example of Extrusion Feature Addition Using an Existing Face as Sketch Reference Plane.
+However, current models and datasets often define all points, lines, and surfaces using coordinates within space, such as DeepCAD.
+This method lacks the dependencies between elements, for instance, a sketch’s reference plane can be defined based on existing points, lines, and surfaces rather than solely relying on the origin and direction vectors.
+This dependency-based approach to defining 3D shapes, which aligns with human modeling methods, enhances robustness, readability, and editability.
+The Figure 1 illustrates an example of adding an extrusion feature to an existing geometry, where the new extrusion is created using an existing face as the reference plane for the sketch.
+Additionally, the vision towers of current multimodal language models, such as OpenAI’s CLIP[24] and Google’s SigLIP [25], tend to focus on macroscopic image analysis, making it difficult to extract fine details, which leads to a significant loss of shape details during generation.
+In this work, we leverage the visual, logical, coding, and general capabilities of visual language models to make the following contributions:
   1. 1.
-    We established OpenECAD dataset.
-Based on existing public CAD datasets, such as SketchGraphs[18], Fusion 360 Gallery[26], and DeepCAD[24], we convert these datasets into ”image-CAD operation code” pairs, including 2D sketch generation and 3D shape generation.
-Using currently available multimodal large language models, such as GPT-4o, we generate high-quality annotated datasets.
-Additionally, we utilize rendering tools to generate various forms of images to enhance the versatility of model inputs.
+    We established OpenECAD datasets.
+We transform the DeepCAD dataset into pairs of “image-CAD operation code”.
+Where feasible, we employ reference planes to establish new sketch planes, moving beyond the sole reliance on points and direction vectors.
+The dimensions within the sketches will be consistent with the absolute space, rather than being scaled as in DeepCAD.
+These changes enhance the editability of the code.
+Additionally, we will utilize rendering tools to generate images including single random view, single isometric and transparent view, and a combination of isometric view and three orthographic views, enhancing the versatility of model inputs.
   2. 2.
-    We trained OpenECAD model.
-Based on existing public small language models and multimodal small language models, such as TinyLLaMA[27], OpenELM[28], Gemma[29], Phi-2, Phi-3-mini[30], and TinyLLaVA[31], we train them using the aforementioned training sets and test their performance in CAD design.
+    We trained OpenECAD models using existing small language models and multimodal small language models, including OpenELM[26], Gemma[27] and Phi-2[28].
+We employed TinyLLaVA’s [29] training methodology, using OpenAI’s CLIP or Google’s SigLIP as vision tower, and training these models with the LLaVA dataset[30] to imbue them with multimodal capabilities.
+Subsequently, we fine-tuned both the language models and vision towers of these architectures with OpenECAD datasets and evaluated their performance in CAD design.
+These models, with the largest containing only 3.1B parameters, demonstrate efficient inference capabilities in single GPU environments.
   3. 3.
-    We integrated the code generated by OpenECAD into CAD tools.
-We design a program that can interface with the APIs of various CAD tools, enabling the generated model code to be directly used in CAD applications.
+    We designed a program that can read the generated CAD code and interface with the APIs of PythonOCC, enabling the reconstruction of 3D models from the code and testing the models.
+Unlike previous models, OpenECAD models require only one or few 2D view(s) to understand 3D shapes, without needing point clouds or Brep files.
+The generated code is highly editable and aligns with human operations in CAD tools, enabling flexible manual modifications.
 2 Related Work
+2.1 Sketch-based 3D Modeling Techniques
+In recent years, researchers have made significant progress in the field of automatic conversion from 2D sketches to 3D models.
+“From sketches to CAM models”[31] proposed a method to directly generate 3D CAM models from 2D sketches by identifying high-level geometric information such as “steps” and “pockets” in 2D line drawings to construct 3D model trees.
+This approach simplifies traditional CAD/CAM workflows, enabling non-professionals to design and manufacture models.
+“Extracting datums to reconstruct CSG models”[32] focused on extracting datums from 2D engineering sketches and reconstructing Constructive Solid Geometry (CSG) models through feature extraction, feature tree construction, and datum definition, effectively translating design intent.
+“Isometric Conversion of Mechanical Sketches”[33] introduced the SFBCM (Sketch Feature-Based Conversion Method), utilizing human perception concepts to convert complex mechanical sketches by progressively detecting and extracting simple features.
+Lastly, “Method to Automatically Convert Sketches of Mechanical Objects”[34] further expanded the types of sketches that can be processed, defining various sketch faces and features, and proposing a detailed algorithmic process that successfully handled sketches of multiple complex mechanical objects.
+These studies collectively advanced the development of sketch-based 3D modeling techniques, demonstrating the feasibility of extracting complex 3D features solely from 2D images.
+Furthermore, they laid a solid theoretical and technical foundation for the subsequent generation of 3D shapes using large neural network models.
 2.2 Small Language and Visual Language Models
@@
 Figure 2: Overview of the OpenECAD Dataset and Model.
-We now present our OpenECAD dataset and model, which includes a new CAD command sequence format designed for language models (see Figure 2).
-The dataset has been improved based on existing CAD datasets and utilizes large language models to generate high-quality natural language descriptions of CAD models.
-We fine-tuned pre-trained vision-language models based on small language models, resulting in the OpenECAD models that are compact in size and capable of running locally, specifically designed for CAD applications.
+We now present our OpenECAD datasets and models (see Figure 2), which include a new CAD command sequence format designed for language models.
+These datasets are converted from the DeepCAD dataset and filtered based on the token count corresponding to the language model, resulting in multiple datasets of varying sizes.
+Additionally, we utilize large language models to generate high-quality natural language descriptions of OpenECAD codes.
+Based on multiple small language models of different sizes, we fine-tuned pre-trained vision-language models, ultimately obtaining several OpenECAD models of varying sizes.
+These models are compact, capable of running locally, and specifically designed for CAD applications.
+The python code generated by the models can produce CAD project files through corresponding API operators of CAD tools.
 4 Generation of Datasets
 In this chapter, we encompass several key components essential for OpenECAD datasets.
-We cover the design and generation of datasets (Section 4.1), the definition of the code format for CAD operation sequences (Section 4.2), and the translation of these sequences (Section 4.3).
-Additionally, we address the generation of natural language descriptions for CAD models (Section 4.4), the production of CAD model view images (Section 4.5), and the overall creation of the CAD dataset (Section 4.6).
-4.1 Design and Generation of Datasets
+We cover the design of datasets (Section 4.1), the definition of the code format for CAD operation sequences (Section 4.2), and the translation of these sequences (Section 4.3).
+Additionally, we address the generation of natural language descriptions for CAD models (Section 4.4), the production of CAD model view images (Section 4.5), and the overall creation of OpenECAD datasets (Section 4.6).
+4.1 Design of OpenECAD Datasets
 CAD models can be represented in various ways.
 At the user interaction level with CAD software, 3D shapes are described by a series of CAD operations to create solid forms.
-For instance, to model a simple table, a user begins by sketching on a 2D plane, drawing a rectangle (a closed curve) and constraining its dimensions.
-After completing the sketch, the user extrudes it to form the tabletop and constrains its thickness.
-Next, the user sketches on the existing model, drawing four squares (four closed curves) and constraining their dimensions.
-After completing this sketch, the user extrudes these shapes to form the table legs, merging them with the previous model and constraining their height.
-This completes the modeling of a simple table (see Figure 3).
-This process constitutes a CAD operation sequence, consisting of two Sketch-Extrusion pairs.
+For example, if we need to add a cuboid to an existing feature, we first need to specify the reference plane for sketching.
+Then, we draw a rectangle on the sketch plane.
+Once the sketch is completed, it forms a closed profile, and the extrude operation is applied to this profile.
+After the extrusion is completed, the addition of the cuboid is finished.
+By repeatedly performing this operation, we can create a complex 3D shape (see Figure 1).
 Clearly, this CAD operation sequence can be described in natural language, as demonstrated above.
 Additionally, the 2D view of a CAD model is the most intuitive way to express it to humans, akin to the flowchart mentioned above.
-Figure 3: A CAD Command Sequence Diagram for Drawing a Table.
 In this work, we aim to generate CAD operation sequences using visual language models.
@@
 To constrain the generation of CAD operation sequences, we use human-understandable inputs such as model images or textual descriptions.
-Therefore, our dataset includes pairs of target model images as input with CAD operation codes and descriptions as output, pairs of target models as input with CAD operation codes as output, and pairs of target model images and descriptions as input with CAD operation codes as output.
-These datasets encompass both 3D CAD models and pure 2D sketches to enhance the model’s ability to generate and constrain 2D sketches.
 4.2 Definition of the Code Format for CAD Operation Sequences
-Mature CAD tools support a rich set of commands, most of which are simplified calls to collections of basic commands—for example, a rectangle is a collection of four line segments and some constraints.
-However, in practice, only a small subset of these commands is widely used.
-Therefore, we currently consider only the most basic and commonly used commands in our dataset (see Table 1).
-Due to the versatility of language models, extending the dataset later is relatively easy.
 Command Name Parameters
@@
 Sketch and its helper
-add_sketchplane origin_point, x_axis, y_axis
+add_sketchplane origin, normal, x_axis
+add_sketchplane_ref extrude, origin, type, (optional values)
 add_profile loops_list
@@
 Table 1: CAD commands and their parameters.
-Sketch.
-Sketch is the core of CAD operations.
-Extrusion is built upon Sketch.
-Curves and constraints form closed loops within a Sketch.
-Multiple closed loops make up the profile of a closed region.
-In our representation, the definition of Sketch begins with defining the sketch reference plane, similar to most CAD tools.
-We then maintain a list of loops and start drawing curves and adding constraints.
-Once a loop is completed, it is stored in the list.
-After all loops are completed, the corresponding profile is generated, and finally, the sketch is created based on the reference plane and the profile.
-The sketch reference plane is defined by the three-dimensional coordinates of the origin and the direction of the sketch’s XY axes.
-Curves.
-In practice, we consider three types of widely used curve commands: drawing a line, drawing an arc, and drawing a circle.
-The line is defined by its starting and ending points, the arc is determined by its starting point, ending point, and midpoint, and the circle is specified by its center and radius.
-Extrusion.
-The extrusion command extrudes a sketch from a 2D plane into a 3D body.
-The extrusion type can be one-sided, symmetric, or two-sided with respect to the sketch plane.
-The command also specifies how to merge the newly extruded 3D body with the previously created shape using one of the operations: creating a new body, or joining, cutting, or intersecting with the existing body.
+Figure 3: Comparison Diagram of Two Definitions for Line, Arc, and Circle.
+Mature CAD tools support a rich set of commands, most of which are simplified calls to collections of basic commands—for example, a rectangle is a collection of four line segments and some constraints.
+However, in practice, only a small subset of these commands is widely used.
+Therefore, we currently consider only the most basic and commonly used commands in our dataset (see Table 1).
+Curve operations include add line, arc, or circle, all of which use the most intuitive variables for definition (as shown in the definition method on the left side of Figure 3).
+Sketch operations include defining the reference plane directly using the origin, direction vector, and x-axis direction, or using existing features to define the reference plane.
+When using existing features, the inputs vary depending on the type, including: using the same reference plane as the reference extrusion, using the top face or the side face obtained from the extrusion.
+When using the side face, the corresponding line must be specified.
+These methods can determine whether to reverse the direction vector or rotate the x-axis.
+Sketch operations also include auxiliary functions such as obtaining the profile and the sketch.
+Extrusion operation requires inputs such as the sketch, operation type, extrusion type, and extrusion length.
+The operation types include simple extrusion, cut, or join with existing features, and the extrusion types include unidirectional or bidirectional extrusion.
+Due to the versatility of language models, extending the dataset later is relatively easy.
 All operation names are chosen to be as meaningful as possible to aid the language model in understanding and using them.
 With these operation commands, we describe a CAD model as a sequence of executable commands in code.
-The pseudo code corresponding to a pair of Sketch-Extrusion operations is shown in the Algorithm 1.
-Algorithm 1 Pseudo Code Example
-$SketchPlane$ = add_sketchplane($origin$, $normal$, $x_{axis}$, $y_{axis}$)
-$Loops$ and $Curves$ are lists.
-$Line1$ = add_line($start$, $end$)
-$Line1$ is appended to $Curves$.
-$Loop$ = add_loop($Curves$)
-$Loop$ is appended to $Loops$.
-$Profile$ = add_profile($Loops$)
-$Sketch$ = add_sketch($SketchPlane$, $profile$, $position$, $size$)
-$Extrude$ = add_extrude($Sketch$, $operation$, $type$, $extent$)
+An example of OpenECAD code is shown in Algorithm 1.
+Algorithm 1 OpenECAD Code Example
+1:  $SketchPlane0$ = add_sketchplane($origin$, $normal$, $x_{axis}$)
+2:  $Loops0$, $Curves0\_0$ = [], []
+3:  $Line0\_0\_0$ = add_line($start$, $end$)
+4:  …
+5:  $Loop0\_0$ = add_loop($Curves0\_0$)
+6:  $Profile0$ = add_profile($Loops0\_0$)
+7:  $Sketch0$ = add_sketch($SketchPlane0$, $Profile0$, $position$, $size$)
+8:  $Extrude0$ = add_extrude($Sketch0$, $operation$, $type$, $extent$)
+9:  $SketchPlane1$ = add_sketchplane_ref($Extrude0$, $origin$, $type$= “line”, $line$= $Line0\_0\_1$)
+10:  …
 4.3 Translation of CAD Operation Sequences
 According to the definitions provided in the previous section, we can convert any CAD model covered by the described operations (in Table 1) into a CAD operation sequence as defined by us.
-However, in practice, different CAD tools or datasets may have slight variations in their definitions of CAD operations (see Figure 4).
-Therefore, we determine and transform each curve-drawing operation according to its degrees of freedom to match the syntax of the OpenECAD dataset.
-We have developed scripts to convert other CAD models or datasets into CAD operation sequence codes for training our VLM.
-Figure 4: Comparison Diagram of Two Definitions for Line, Arc, and Circle.
+However, in practice, different CAD tools or datasets may have slight variations in their definitions of CAD operations.
+For curve operations, we determine and transform each curve-drawing operation according to its degrees of freedom to match the syntax of the OpenECAD dataset.
+For the DeepCAD dataset, we do not need to convert their definition methods; we only need to convert them into OpenECAD codes to train our vision language model (VLM).
+However, for other datasets such as AutoDesk Fusion, we need to design scripts to calculate the corresponding relationships based on geometry (see Figure 3).
+Algorithm 2 Find reference plane
+1:  $t$ is target plane’s normal vector
+2:  for all existed extrudes: $E$ do
+3:     $n$ is $E$’s normal vector
+4:     if $n$ is parallel with $t$ then
+5:        if two plane’s distance is 0 or extent_size of $E$ then
+6:           reference plane found.
+7:           return
+8:        else
+9:           $E$ can’t contain reference plane.
+10:        end if
+11:     else if $n$ is perpendicular with $t$ then
+12:        for all lines in sketch of $E$: $l$ do
+13:           if $l$ on the target plane then
+14:              reference plane found.
+15:              return
+16:           end if
+17:        end for
+18:     else
+19:        $E$ doesn’t contain reference plane.
+20:     end if
+21:  end for
+For sketch operations, the primary focus is on defining the sketch reference plane, and the functions that assist in constructing the sketch do not require special conversion operations.
+For the DeepCAD dataset, for planes that can only be defined directly using the origin and vectors, we do not need to convert their definition methods, as the direction vector and x-axis can be easily calculated.
+We only need to convert them into OpenECAD codes.
+However, for reference planes that can be defined using existing features, we need to design an algorithm to find the appropriate plane.
+We will traverse all previously defined sketch planes and the planes obtained from extrusions (limited to flat surfaces, excluding curved surfaces) to find a suitable reference plane.
+This algorithm is shown in Algorithm 2.
+For extrusion operations, no special conversion is required.
 4.4 Generation of Natural Language Descriptions for CAD Models
@@
 In this section, we will briefly explain how to generate image materials for VLM training, specifically by creating 2D views of CAD models.
-We first use the PythonOCC[33] tool to write scripts that generate step files (B-Rep format) based on the CAD operation sequence codes.
-Then, we render the step files using PythonOCC.
-Finally, we select a direction that forms a random angle with the isometric view direction vector to generate 2D views as image materials.
+We first use the PythonOCC tool to write scripts that generate step files (B-Rep format) based on the CAD operation sequence codes.
+Then, we generate three types of 3D shape’s images: Default View, Transparent View, and Orthographic Views.
+The Default View is a 2D view of the 3D shape generated by selecting a direction that forms a random angle with the isometric view direction vector.
+The Transparent View makes the 3D shape transparent to reveal features that are not visible from a certain direction.
+The Orthographic Views consist of a single image composed of the front view, top view, left view, and isometric view.
+We use PythonOCC to render these step files and generate the corresponding images.
 4.6 Creation of CAD Dataset
-Based on the above method, we generated a 3D CAD dataset derived from DeepCAD dataset[24] and Fusion 360 Gallery Reconstruction Dataset[26].
-We excluded designs that were generated incorrectly or had overly long CAD command sequences due to the number of tokens that small language models can accept is limited, resulting in a total of 130,239 3D designs.
-The statistical distribution of the number of ”Sketch-Extrusion” steps in the 3D dataset is shown in the Figure 5.
-Figure 5: The statistical distribution of the number of ”Sketch-Extrusion” steps in OpenECAD 3D dataset.
+Based on the above method, we generated 3 OpenECAD datasets derived from the DeepCAD dataset.
+They contain 100,000, 150,000, and 200,000 “image-code” pairs, respectively, with code lengths limited to 1024 tokens, 2048 tokens, and 3072 tokens.
+We excluded designs that were generated incorrectly or had overly long CAD command sequences due to the limited number of tokens that small language models can accept, as well as designs used for the validation set.
+The distribution of “Default View”, “Transparent View”, and “Orthographic Views” in the datasets is shown in the Table 2.
+“$x$_directout_$y$” indicates that the dataset contains $x$ “image-code” pairs, with the code length limited to $y$ tokens.
+This dataset is used for directly outputting complete OpenECAD code from image input.
+Dataset Name Default View Transparent View Orthographic Views
+100k_directout_1k 62500 25000 12500
+150k_directout_2k 90000 37500 22500
+200k_directout_3k 120000 50000 30000
+Table 2: The distribution of OpenECAD datasets.
+The statistical distribution of the number of “Sketch-Extrusion” steps in the OpenECAD datasets is shown in the Figure 4.
+As shown in the figure, designs with only one “sketch-extrude” pair still constitute the majority of datasets, especially in datasets with strict token limits.
+This is partly because the token limit prevents OpenECAD code from being long enough to accommodate more “sketch-extrude” pairs in a single design, and partly because such designs make up the majority of the DeepCAD dataset.
+In datasets where the token limit is less restrictive, more complex designs have a more significant proportion.
+Figure 4: The statistical distribution of the number of “Sketch-Extrusion” steps in OpenECAD datasets.
 5 Training and Fine-tuning Vision Language Models
@@
 5.1 The choice of language models and multimodal approach
-Due to the high confidentiality requirements of manufacturing enterprises and the substantial CPU and GPU resources typically occupied by CAD design tools, we chose OpenELM[28] (450M), Gemma[29] (2B), Phi-2 (2.7B) and Phi-3 mini[30] (3.8B) with relatively small parameter sizes as the foundation for our language model.
-OpenELM is an efficient language model family developed by Apple.
-It employs a layer-wise scaling strategy to allocate parameters efficiently within each layer of the transformer model.
-This approach enhances accuracy while maintaining computational efficiency.
-Gemma is a family of lightweight, text-to-text, decoder-only large language models developed by Google.
-Phi, developed by Microsoft Research, is another series of advanced small language models (SLMs).
-This series includes Phi-2 (2.7B), Phi-3-mini (3.8B), Phi-3-small (7B), and Phi-3-medium (14B), each with different parameter sizes.
-Despite their compact parameter scales, Phi-3 models achieve language understanding and reasoning capabilities comparable to larger models.
+Due to the high confidentiality requirements of manufacturing enterprises and the substantial CPU and GPU resources typically occupied by CAD design tools, we chose OpenELM[26] (450M), Gemma[27] (2B) and Phi-2 (2.7B)[28] with relatively small parameter sizes as the foundation for our language model.
 These models are trained on high-quality data and optimized algorithms.
 It is worth noting that the context acceptance length for OpenELM (450M) is 2048 tokens.
-In contrast, Gemmma and Phi has a minimum context acceptance length of 3072 tokens, with an option to use a version supporting up to 128k tokens at the cost of some content accuracy.
-To endow the language model foundation with multimodal capabilities, we selected OpenAI’s CLIP[34] and Google’s SigLIP [35] as visual encoders for OpenELM (450M), Gemma (2B) and Phi-2 (2.7B).
+In contrast, Gemma and Phi has a minimum context acceptance length of 3072 tokens, with an option to use a version supporting up to 128k tokens at the cost of some content accuracy.
+To endow the language model foundation with multimodal capabilities, OpenAI’s CLIP[24] and Google’s SigLIP are selected [25] as visual encoders for OpenELM (450M), Gemma (2B) and Phi-2 (2.7B).
 These models are trained on large datasets consisting of images paired with corresponding text descriptions, combining visual and textual information to make predictions.
-To train the entire model with multimodal conversational abilities, we employed the LLaVA[36] method and its dataset.
+To train the entire model with multimodal conversational abilities, the LLaVA[30] method and its dataset are employed.
 LLaVA combines a visual encoder and a large language model in an innovative multimodal approach, enabling comprehensive visual and language understanding.
 LLaVA uses GPT-4 to generate multimodal language-image instruction data.
-For training the model, we utilized the training framework provided by TinyLLaVA[31].
-For Phi-3 mini, since Microsoft has already released its multimodal version, Phi-3 Vision (4.2B), which was trained using methods similar to LLaVA, we will directly use the pre-trained model provided by Microsoft.
+We utilized the training framework and the pretrained models provided by TinyLLaVA[29].
 5.2 Fine-tuning on a pre-trained visual language model
-OpenECAD 0.55B & 0.89B
-CPU Intel Core i9-7960X
-GPU Nvidia GeForce GTX 1080 Ti
-RAM DDR4-2666 32 GB
+OpenECAD 0.55B & 0.89B & 3.1B
+CPU Intel Xeon Platinum 8352V
+GPU 2 x Nvidia GeForce RTX 4090
+RAM 90 GB
 OS Ubuntu Server 22.04 LTS
-OpenECAD 2.4B & 3.1B
-CPU AMD EPYC 9754 Processor (18 vCPU)
-GPU 2 x Nvidia GeForce RTX 4090D
-RAM 60 GB
-OS Ubuntu Server 22.04 LTS
-OpenECAD 4.2B
-CPU AMD EPYC 9754 Processor (18 vCPU)
-GPU Nvidia GeForce RTX 4090D
-RAM 60 GB
+OpenECAD 2.4B
+CPU Intel Xeon Platinum 8352V
+GPU 2 x Nvidia GeForce RTX 4080 Super (as one vGPU)
+RAM 90 GB
 OS Ubuntu Server 22.04 LTS
@@
 This significantly reduces the number of trainable parameters for downstream tasks.
-We performed LoRA fine-tuning on the pre-trained OpenELM-CLIP (0.55B) and OpenELM-SigLIP (0.89B) models using an Nvidia GeForce GTX 1080 Ti (11 GB VRAM).
-The Rank and Alpha for LoRA were set to 128 and 256, respectively.
-The learning rate is $10^{-4}$.
-Training for one epoch with our 3D dataset took approximately 60 hours.
-Through fine-tuning, we obtained the OpenECAD 0.55B and 0.89B models.
-Similarly, we used 2 Nvidia GeForce RTX 4090 D (24 GB VRAM) to perform LoRA fine-tuning on the Gemma-SigLIP (2.4B) and Phi-2-SigLIP (3.1B) model.
-The Rank and Alpha for LoRA were set to 128 and 256, respectively.
-The learning rate is $10^{-4}$.
-Training for one epoch with our 3D dataset took approximately 14 hours.
-Through fine-tuning, we obtained the OpenECAD 2.4B and 3.1B models.
-We used an Nvidia GeForce RTX 4090 D (24 GB VRAM) to perform LoRA fine-tuning on the Phi-3 Vision (4.2B) model.
-The LoRA rank and alpha were set to 256 and 512, respectively.
-The learning rate is $10^{-4}$.
-Training with our dataset for one epoch took approximately 20 hours.
-Through fine-tuning, we obtained OpenECAD 4.2B model.
 The specific development, training, and runtime environments are shown in the Table 3.
-The loss curves for the OpenECAD 0.55B, 0.89B, 2.4B, 3.2B and 4.2B models during training are illustrated in the Figure 6.
-Due to the different batch sizes used during training, the loss curves are proportionally scaled.
-As shown, the models converge after a single round of fine-tuning.
-Figure 6: The loss curves for the OpenECAD 0.55B, 0.89B, 2.4B, 3.2B and 4.2B models during training.
+The loss function used was CrossEntropyLoss, and the training method was instruction fine-tuning.
+The learning rate was set to $10^{-4}$, and the learning rate scheduler type was cosine.
+The training of all OpenECAD models were conducted with a train batch size of 2 per device and gradient accumulation steps of 2.
+The Rank and Alpha for LoRA were set to 128 and 256.
+Besides using LLM to add comments to a small portion of OpenECAD datasets’ codes for data augmentation, further data augmentation is of limited significance since the datasets are already sufficiently large for fine-tuning.
+We performed LoRA fine-tuning on the pre-trained OpenELM-CLIP (0.55B), OpenELM-SigLIP (0.89B) and Phi-2-SigLIP (3.1B) models.
+The training was conducted on 2 Nvidia GeForce RTX 4090 GPUs.
+Under the environment described in Table 3, fine-tuning on the pre-trained OpenELM-CLIP (0.55B) and OpenELM-SigLIP (0.89B) models with the OpenECAD 100k_directout_1k dataset took approximately 2 hours, fine-tuning on the pre-trained Phi-2-SigLIP (3.1B) model with the OpenECAD 200k_directout_3k dataset took approximately 20 hours.
+Similarly, we used 2 Nvidia GeForce RTX 4080 Super GPUs as a single vGPU to perform LoRA fine-tuning on the Gemma-SigLIP (2.4B) model.
+Under the environment described in Table 3, fine-tuning on the pre-trained Gemma-SigLIP (2.4B) model with the OpenECAD 150k_directout_2k dataset took approximately 30 hours
+The loss curves of the first 1500 steps for the OpenECAD 0.55B, 0.89B, 2.4B, and 3.1B models during training are illustrated in Figure 5.
+As shown, the models converge after the first 1500 steps of fine-tuning.
+Figure 5: The partial loss curves for the OpenECAD 0.55B, 0.89B, 2.4B, and 3.1B models during training.
 6 Experiments
@@
 Following this, we test the models, obtain the results (Section 6.3), and analyze them (Section 6.4).
+We also provide an example of drawing a simple table (Section 6.5) and discuss future work (Section 6.6).
 6.1 Evaluation Metrics
@@
 Scoring Item Score (out of 100)
-Is the code executable? 10
-Accuracy of 3D shape construction 90
-For each Loop-Extrusion pair (out of 100)
-- 2D sketch positioning 20
-- Accuracy of 2D Loop 60
-- 3D operation 20
+Is the code executable? 10 * exec_flag
+Accuracy of All Curves 45 * acc
+Accuracy of Loops count 5 * acc
+Loops integrated score 40 * score
+Fo each Loop’s score (out of 100)
+If loop is absolutely correct 100
+Else calculate acc of curves 90 * acc
 Table 4: Scoring Table.
@@
 To assess the generative capability of our model, we designed a scoring algorithm.
-First, the algorithm verifies if the generated code can be executed directly by the API operator to create a model.
-If it cannot run, it scores 0 points.
-If it can run, we then use a visual language model (such as GPT-4o) to compare the generated model’s view with the target input, evaluating each basic feature.
-Each basic feature corresponds to a 2D loop.
+The algorithm verifies if the generated code can be executed directly by the API operator to create a model, the accuracy of curve operations, the accuracy of the number of loops, and provides a comprehensive score for the loops.
 We separately evaluate the loops in the 2D sketches during the modeling process to ensure precise assessment, as these loops may or may not be considered a single sketch during modeling.
 We score the results based on executability and the correctness of the 3D shapes (see Table 4).
-Each Loop-Extrusion pair is considered a basic 3D shape.
-If the number of generated Loop-Extrusion pairs is insufficient, points will be deducted proportionally.
-For each Loop-Extrusion pair, we score based on the reference plane position of the 2D sketch, the extrusion operation, and the drawing of the 2D loop.
+For the comprehensive scoring of loops, we first check if the two loops are completely identical, meaning the types and order of the curves are exactly the same, and assign a score of 100.
+If not, the score is recorded as 90% of the curve accuracy.
 The specific calculation method is shown in the following formula.
-$score=10R+\frac{90}{N_{s}}\sum_{i=1}^{N_{s}}(20P_{i}+20O_{i}+60L_{j})$ (1)
-where $R$ denotes executability, $N_{s}$ represents the number of Loop-Extrusion pairs, $P_{i}$ and $O_{i}$ indicate the correctness of the 2D sketch position and 3D extrusion operation for the $i$-th Sketch-Extrusion pair, and $L_{j}$ denotes the accuracy of the curves within the loop.
-$R,P_{i},O_{i},L_{j}\in[0,1]$.
+$score=10e+45acc_{c}+5acc_{l}+\frac{40}{100L}\sum_{i=1}^{L}(10s_{i}+90acc_{i})$ (1)
+where $e$ denotes executability, $acc_{c}$ represents accuracy of curve operations, $acc_{l}$ represents accuracy of the number of loops, $L$ represents the number of Loops, $s_{i}$ and $acc_{i}$ indicate whether the loop is absolutely correct and the accuracy of curve operations for the $i$-th loop.
+$e,acc_{c},acc_{l},s_{i},acc_{i}\in[0,1]$.
 6.2 CAD Designs for Test
 To test the model, we require high-quality CAD designs.
-We extracted approximately 100 3D designs from the training set, maintaining a ratio of 10:5:3:2 for designs containing 1, 2, 3, and 4 or more Sketch-Extrusion pairs, respectively.
-After filtering and removing duplicates, we obtained 57 unique 3D designs.
-These designs were redrawn using CAD tools such as AutoDesk Fusion, and appropriate views were selected to generate the input images.
+We extracted 625 3D designs from the OpenECAD dataset that were not included in the training set.
+Similar to the method used for generating the training set, we created their standard views, perspective views, and four-angle views.
 Some examples are shown in the Figure 6.
-The resolution of the images is 640x400.
+The resolution of the images is 640x480.
+The statistical distribution of the number of “Sketch-Extrusion” steps in the is shown in the Figure 7.
+Figure 7: The statistical distribution of the number of “Sketch-Extrusion” steps of CAD designs for evaluating.
 6.3 Evaluation Results
-Figure 8: Comparison of some successful rendered test results illustrations.
 6.3.1 The Inference Speed of the Models
-The input was standardized to a 640x400 image of a 3D shape, with the prompt:
-”This image is a view of a 3D model from a certain angle. Please try to use Python-style APIs to render this model.”
-For OpenECAD 0.55B and 0.89B, we used an Nvidia GeForce GTX 1080 Ti (11 GB VRAM) for inference, with a peak memory usage of less than 5GB.
-Getting a 3D shape’s code took approximately 20 seconds.
-For OpenECAD 2.4B, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) for inference, with a peak memory usage of less than 9GB.
-For OpenECAD 3.2B and 4.2B, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) for inference, with a peak memory usage of less than 18GB.
-Getting a 3D shape’s code took approximately 45 seconds.
+The input was standardized to a 640x480 image of a 3D shape, with the prompt:
+“This image is a(4) view(s) of a 3D model from a certain angle. Please try to use Python-style APIs to render this model.”
+For all OpenECAD models, we used an Nvidia GeForce GTX 1080 Ti (11 GB VRAM) for inference.
+OpenECAD 0.55B and 0.89B models cost a peak memory usage of less than 4GB.
+Getting a 3D shape’s code took approximately 30 seconds.
+For OpenECAD 2.4B, it infers with a peak memory usage of less than 9GB.
+OpenECAD 3.1B infers with a peak memory usage of less than 11GB.
+Getting a 3D shape’s code took approximately 45 seconds using OpenECAD 2.4B and 3.1B models.
 6.3.2 Generating Code and Rendering
-We tested OpenECAD versions 0.55B, 0.89B, 2.4B, 3.2B and 4.2B, and also tested GPT-4o.
+Figure 8: Some successful rendered test results illustrations
+We tested OpenECAD versions 0.55B, 0.89B, 2.4B and 3.1B, and also tested GPT-4o-mini (2024-07-18 version).
 Some successfully rendered test results are illustrated in Figure 8.
-Since GPT-4o was not trained specifically for the use case described in this paper, we used the following prompt for testing:
-”I am going to provide you with a 3D shape’s corresponding Python code for its modeling process.
+Since GPT-4o-mini was not trained specifically for the use case described in this paper, we used the following prompt for testing:
+“I am going to provide you with a 3D shape’s corresponding Python code for its modeling process.
 After that, I will input images of other 3D shapes, and please provide the corresponding Python code for them.
@@
 Please write the codes of it in the format above.”
-During the testing of OpenECAD, the maximum number of new tokens that 0.55B, 0.89B, 3.2B and 4.2B models can generate is 1536, 1024, 2048, 3072 and 3072, respectively.
+During the testing of OpenECAD, the maximum number of new tokens that the 0.55B, 0.89B, 2.4B, and 3.1B models can generate is 1536, 1152, 2048, and 2048 respectively.
 Due to these context limitations, there are instances where the generated code is incomplete.
-GPT-4o, due to its support for a longer context, did not encounter issues with incomplete code generation.
+GPT-4o-mini, due to its support for a longer context, did not encounter issues with incomplete code generation.
 Similar to generating datasets, we use the PythonOCC tool to write a library that attempts to directly run the generated code to create CAD projects and output them as STEP files (B-Rep format).
@@
 Since the generated programs from the model cannot ensure complete accuracy, errors may occur that the API operator cannot handle, leading to model rendering failures.
-Based on the scoring method mentioned above, we obtained the scores, the number of executable results, the number of incomplete results, the number of runtime errors, and the number of completely correct results for each model, as shown in the Table 4.
-Model Overall Executable Unfinished Run with Completely
-Score Errors Correct
-0.55B 46.52 36 21 0 15
-0.89B 37.19 22 35 0 19
-2.4B 59.86 50 7 0 20
-3.1B 55.39 53 0 4 16
-4.2B 51.91 51 4 2 14
-GPT-4o 15.44 52 0 5 4
+Based on the scoring method mentioned above, we obtained Curves accuracy, Loops count accuracy, Loops score, Execution score, and Overall score for each model and input type, as shown in the Table 5.
+Additionally, we calculated the scores excluding designs with only one “Sketch-Extrusion” pair.
+Default View as input Default View without only 1 pair designs
+Model 0.55B 0.89B 2.4B 3.1B gpt-4o-mini 0.55B 0.89B 2.4B 3.1B gpt-4o-mini
+Curves accuracy 85.95 88.97 91.23 87.71 40.11 68.65 74.27 76.01 70.29 25.81
+Loops accuracy 84.27 86.79 90.62 86.98 75.99 74.23 76.72 76.43 70.63 57.10
+Loops score 77.73 80.03 84.20 81.46 34.45 65.50 68.57 72.96 69.18 25.07
+Execution score 99.04 96.64 95.36 94.40 40.96 98.92 90.32 85.48 84.95 24.73
+Overall score 83.89 86.05 88.80 85.84 39.72 70.70 73.72 75.76 71.33 26.97
+Transparent View as input Transparent View without only 1 pair designs
+Model 0.55B 0.89B 2.4B 3.1B gpt-4o-mini 0.55B 0.89B 2.4B 3.1B gpt-4o-mini
+Curves accuracy 84.57 88.03 90.68 86.72 47.38 70.00 73.65 78.52 70.78 30.00
+Loops accuracy 83.12 84.98 88.74 85.90 76.20 75.80 76.49 76.57 73.36 58.54
+Loops score 77.57 79.83 83.94 81.32 41.02 66.68 68.71 75.41 71.71 28.88
+Execution score 99.20 97.76 95.04 93.44 44.48 99.46 93.55 85.48 82.26 25.81
+Overall score 83.16 85.57 88.32 85.19 45.99 71.91 73.81 77.88 72.43 30.56
+Orthographic View as input Orthographic View without only 1 pair designs
+Model 0.55B 0.89B 2.4B 3.1B gpt-4o-mini 0.55B 0.89B 2.4B 3.1B gpt-4o-mini
+Curves accuracy 85.28 87.02 90.34 87.71 31.78 67.63 72.08 74.91 69.32 22.96
+Loops accuracy 82.73 84.46 89.17 87.02 71.11 70.31 75.17 73.89 70.08 54.86
+Loops score 76.50 79.02 83.21 81.14 26.75 63.18 68.47 72.15 68.50 21.97
+Execution score 99.68 97.44 95.52 95.20 34.88 98.92 91.94 86.56 84.41 23.12
+Overall score 83.08 84.73 87.95 85.80 32.04 69.11 72.78 74.92 70.54 24.17
 Table 5: Evaluation Results Table.
-As shown in Table 4, OpenECAD 2.4B achieved the highest overall score and the most fully correct outputs.
-Both 0.55B and 0.89B models produced a significant amount of incomplete code.
-The 0.89B model’s higher token consumption for image inputs reduced the number of new tokens it could generate, leading to more incomplete outputs compared to the 0.55B model, and thus a lower overall score.
-The 3.1B and 4.2B models achieved relatively high scores, but the number of fully correct generations was lower than that of the 0.55B and 0.89B models.
-All OpenECAD models demonstrated significant advantages over the unrefined GPT-4o, whose sole advantage was its ability to handle the entire code due to its extended context length.
+As shown in Table 5, OpenECAD 2.4B achieved the highest overall score.
+The 0.55B and 0.89B models, despite generating the fewest tokens, scored relatively high in executability.
+This might be because their inability to understand complex models leads them to finish code generation prematurely.
+The 3.1B model did not perform as expected, possibly due to the Phi series models having inferior geometric understanding compared to the Gemma series models.
+All OpenECAD models demonstrated significant advantages over the unrefined GPT-4o-mini, whose sole advantage was its ability to handle the entire code due to its extended context length.
+From the perspective of using different types of view as input, using single-view transparent images as input shows a significant advantage in designs with more than one “sketch extrusion” pair because they contain more information.
+However, using four views as input (isometric view + three orthographic views) performs worse, possibly because the vision tower cannot handle so much information, or the language model cannot establish connections between them.
+The additional views increase the amount of irrelevant information, adding pressure to the model.
 6.4 Analysis of OpenECAD’s outputs
@@
 6.4.2 Generation Error Analysis
-Figure 9: Typical errors in generated results.
-In the generated results, several typical errors were observed (see Figure 9), including:
-  • 1.
-    Case 1 and Case 2: For complex and irregular shapes, the corresponding sketches often fail to generate correctly, or the profiles of the 2D sketches are inaccurately created.
+In the generated results, several typical errors were observed in the Table 6.
+Error Case (Left: target, Right: generation result) Analysis
+[Uncaptioned image] For complex and irregular shapes, the corresponding sketches often fail to generate correctly, or the profiles of the 2D sketches are inaccurately created.
 Possible reasons include the vision tower’s low resolution, which hinders accurate recognition of complex shapes, and the language model’s inability to comprehend complex shapes from the CAD operation code.
 These issues ultimately result in incorrect loops or intersecting loops, leading to erroneous sketch profiles and broken extrusions.
-  • 2.
-    Case 3 and Case 7: The model lacks spatial reasoning capabilities, resulting in incorrect handling of the size proportions between sketches and extrusions.
-This leads to extrusions that are either too long or too small, causing discrepancies in size relative to other parts.
-  • 3.
-    Case 4: When multiple ”Sketch-Extrusion” steps are involved, the positions of the sketches in subsequent steps are often incorrect, leading to errors in the final overall model.
-  • 4.
-    Case 5: Due to the minimal color difference between the outer surface of a solid cylinder and the inner wall of a hollow cylinder, the primary distinguishing feature is the inner wall circle at the top.
+[Uncaptioned image] The model lacks spatial reasoning capabilities, resulting in incorrect handling of the size proportions between sketches and extrusions.
+This leads to extrusions that are either too long or too small, or sketches that are either too large or too small, causing discrepancies in size relative to other parts.
+The lack of numerical understanding is a common issue with language models, which can be relatively easily resolved by post-editing the extent size of the extrusions or the size of curves in the sketches in the CAD code or project.
+[Uncaptioned image] When multiple “Sketch-Extrusion” steps are involved, the positions of the sketches in subsequent steps are often incorrect, leading to errors in the final overall model.
+This can be resolved by post-editing the reference or absolute positions of the sketches in the CAD code or project.
+However, for complex models, this solution might be more challenging.
+[Uncaptioned image] Due to the minimal color difference between the outer surface of a solid cylinder and the inner wall of a hollow cylinder, the primary distinguishing feature is the inner wall circle at the top.
 This makes it difficult for the model to differentiate between solid and hollow cylinders, leading to confusion.
 This issue also extends to other similar shapes with holes.
-  • 5.
-    Case 6: The difficulty in distinguishing concave and convex shapes results in erroneous models.
+[Uncaptioned image] The difficulty in distinguishing concave and convex shapes results in erroneous models.
 This issue may arise because the vision tower can detect these shapes but cannot differentiate between them, or because the language model fails to generate the correct code, leading to sketch errors.
-  • 6.
-    Case 8: Missing part(s). For some relatively subtle features, the model struggles to accurately identify them, resulting in missing features.
-Based on the above results and typical errors, we can infer the current shortcomings of the OpenECAD model:
-  • 1.
+Table 6: Error Cases and Related Analysis
+Based on the results and typical errors mentioned in the Table 6, we can infer the current shortcomings of the OpenECAD model:
   • The model has weak handling capabilities for complex sketches.
 This includes issues such as the vision tower’s low resolution, which leads to an inability to recognize shapes, and the language model’s failure to correctly process complex shapes, resulting in erroneous sketch generation.
-  • 2.
   • The generated models have poor size accuracy, particularly in handling the proportions between sketch dimensions and extrusion sizes.
 This results in mismatched parts, as the model fails to ensure consistent sizes between the sketches and extrusions of different components.
-  • 3.
-    For complex shapes involving multiple ”Sketch-Extrusion” steps, the model struggles to generate accurate results.
+  • For complex shapes involving multiple “Sketch-Extrusion” steps, the model struggles to generate accurate results.
 It has difficulty remembering the previously drawn parts through the context, leading to errors.
-  • 4.
   • Due to the current work focusing only on basic instructions without a combined instruction set like drawing rectangles in one command, the CAD operation codes for 3D shapes tend to be lengthy.
 Moreover, due to contextual limitations of the model, especially with OpenECAD 0.55B and 0.89B using OpenELM which supports only 2048 tokens, and considering that image inputs also consume tokens, there are situations where the code cannot be completed.
+6.5 An Example of Using OpenECAD models
+Figure 9: An example of using OpenECAD 2.4B model to draw a table.
+In this section, we will attempt to use the OpenECAD model by inputting a hand-drawn simple table (see Figure 9), generating the corresponding OpenECAD code, changing the length of 4 legs of the table and then converting it into a step file using the PythonOCC API.
+We input a hand-drawn table’s photo to the OpenECAD 2.4B model, with the prompt:
+“This image is a hand-drawn view of a 3D model from a certain angle. Please try to use Python-style APIs to render this model.”
+The model should respons:
+“Of course, here are the codes:$<$OpenECAD code here$>$
+Algorithm 3 Some OpenECAD Code of the Table
+1:  SketchPlane0 = add_sketchplane(origin= [0., 0., 0.], normal= [0., 0., 1.], x_axis= [1., 0., 0.]))
+2:  Loops0, Curves0_0 = [], []
+3:  Line0_0_0 = add_line(start= [-1000., -750.], end= [1000., -750.])
+4:  …
+5:  SketchPlane1 = add_sketchplane_ref( Extrude0, origin = [0., 0.], type = “sameplan”, reverse = True)
+6:  …
+7:  Extrude1 = add_extrude(sketch= Sketch1, operation= 1, type= 0, extent_one= 1000., extent_two= 0.)
+8:  {extent_one is the length of one of 4 legs of the table.}
+Some key code snippets are shown in the Algorithm 3.
+We changed the extent_one of the table legs extrusion from 1000 to 500.
+We converted the code before and after the change into STEP files using a conversion tool and opened them with a CAD tool.
+As shown in the Figure 9, the table legs are indeed shortened by half.
+Additionally, besides directly modifying the OpenECAD code, it is also possible to convert the OpenECAD code into project files for other 3D CAD tools.
+For instance, AutoDesk Fusion supports all operations in OpenECAD.
+We can use tools and the AutoDesk Fusion API to convert OpenECAD code into AutoDesk Fusion project files.
+This way, we can fully utilize all the features of AutoDesk Fusion to continue modifying the design.
+Due to the limitations of the model’s context length, it is currently challenging for the model to directly modify the generated OpenECAD code based on requirements through dialogue, as the context length does not support accommodating two complete OpenECAD code segments.
 6.6 Future Work
@@
 Based on the above analysis, our future work will focus on the following areas:
-  • 1.
   • Enhancing the dataset involves several key strategies: expanding support for additional CAD instructions, annotating critical shape dimensions in input images to aid accurate size-based shape drawing by the model, selecting optimal angles for input images to showcase all features comprehensively, and employing diverse rendering tools to enhance the model’s ability to recognize shapes under varied lighting conditions and artistic styles.
-  • 2.
   • Improving the visual model is crucial as the current resolution limits its ability to distinguish between certain shapes.
@@
 Additionally, integrating an image segmentation module like FastSAM[41] before the visual model can aid in identifying which parts compose the 3D shapes.
-  • 3.
   • Improving the language model’s capabilities is essential, as the current model supports a limited context length and has weak code generation abilities, resulting in incomplete or erroneous code.
@@
 Additionally, implementing Retrieval-Augmented Generation (RAG) [42] and prompting techniques can further enhance the model’s contextual understanding and code generation capabilities.
-  • 4.
   • Enhancing feedback integration with CAD tools.
@@
 The rendering results can then be immediately returned to the model, helping it understand the current drawing status, identify potential errors, correct them, and continue generating code based on the target shape and the already rendered results.
-At the same time, we have added a Constraints section to the dataset.
-We consider nine types of widely used constraint commands (see Table 5): making a line horizontal or vertical, fixing the size of a curve, making two points of curves coincident, making two lines parallel, perpendicular, or at a fixed angle, and making two curves tangent or mirrored.
-These constraints reduce the degrees of freedom in the sketch.
-A sketch with exactly zero degrees of freedom is fully defined, while one with fewer than zero degrees of freedom is over-constrained.
-We generated a 2D CAD dataset derived from the SketchGraphs dataset[18], resulting in a total of 628,499 2D sketches.
-The statistical distribution of the number of curves and constraints operations in the 2D dataset is shown in the Figure 10.
-Command Name Parameters
-Constraints Series
-make_horizontal(vertical) line0
-fix_size curve0, size
-make_coincident point0, point1
-make_parallel line0, line1
-make_perpendicular line0, line1
-make_tangent curve0, curve1
-make_mirror curve0, curve1 (same type)
-make_angle line0, line1, clockwise
-Table 5: CAD commands and their parameters.
-Figure 10: The statistical distribution of the number of curves and constraints operations in OpenECAD 2D dataset.
 6.6.2 Applications
-Figure 11: An example of a real-world application of the model.
 Currently, OpenECAD primarily serves as a simple CAD operation generation model, focusing on generating models from images.
-However, by improving the dataset, we can train models using similar methods and apply them to downstream applications (like Figure 11).
+However, by improving the dataset, we can train models using similar methods and apply them to downstream applications.
 These applications include, but are not limited to:
-  • 1.
   • Assisting users in operating CAD design tools by guiding them on how to proceed with CAD operations to build or modify models based on their requirements and existing CAD actions.
-  • 2.
   • Enhancing the ability to handle user CAD needs, such as design reusability and designing connectors for existing parts.
-  • 3.
   • Introducing CAE-related knowledge to the model to address some CAE issues, such as structural reinforcement, material reduction, and kinematic simulation, during the CAD phase.
@@
 OpenECAD aims to address the challenge of CAD model generation using visual language models and has successfully generated some relatively simple 3D shapes.
-We offer 5 sizes of OpenECAD models, ranging from 0.55B to 4.2B parameters.
+We offer 4 sizes of OpenECAD models, ranging from 0.55B to 3.1B parameters.
 CRediT authorship contribution statement